A Complete Bibliography of Publications in the ICES Journal of Marine Science (1950–1959)

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

 E-mail: beebe@math.utah.edu, beebe@acm.org, beebe@computer.org (Internet)
            WWW URL: https://www.math.utah.edu/~beebe/

05 October 2023 Version 1.03

Title word cross-reference

14 [Nie52, Nie59, RV54].

0-group [Woo59].

10th [Kor50]. 1949 [Kor50]. 1955 [Ano57e].

Åge [Fri59a]. Aberdeen [Par51]. Abundance [KJ51]. acarne [Cou52b]. Accumulated [Tai51]. Accuracy [Her51, Whi57]. Accurate [CM53]. Acid [Far58]. Acoustical [Wes53a]. Action [AE58, Bag58, WP50]. active [Nie52]. Ad [Ano57e]. Adjacent [Wyr56]. Affected [BGM59]. Affecting [DD50, Dee54, Whi57, Woo57]. Africa [Lon59]. African [Let52]. Age [Fro50, FK59, Men50, Mos54]. Agency [Nie59]. Air [BD57, Sch52]. alalunga [Dol52]. Albacore [Fra53]. Alfred [Büc54]. Algae [Men59]. Amino [Far58]. Ammodytes [For50]. Ammodytidés [For50]. Ammonia [Atk57, Føy50, SA59]. amongst [PS59]. Analyses [Rob50]. Analysis [Bag58, Bur58, Ein52, Jon59]. Anchois [CF50]. Anchovies [CF50]. Angelfish [Men59]. Anglian [Bur58, CB59, Hod58]. Anguilla [Dee52, Dee58]. Angular [AP58]. Animal [Bla56, Dol56]. Animaux [Dol56]. Apparatus [Cox54a, Nie50, Bos59, BR58]. Appareillage [BR58]. Apparent [Bla56]. Appearance [MR58]. Applicable [BW58]. Application [Cab50, Mar54a, Mar57, Tai51]. Archipelago [Sim51a]. Arctic [Mar58, Sma59]. Arcto [Cor58]. Arcto-Norwegian [Cor58]. Areas [Sch52]. Artificial [Dra58, Wal56]. Aspects [Jac54, Wes53a]. Associated [Bag58, Gib59a]. Atlantic [Ano54d, Dol56, Mau52, Nik58, Pos54, Ras59]. Atlantique [Ano54d, Dol56, Pos54]. Atlantiques [Mau52]. Atlas [Hod51]. Atomic [Fon56]. Atomique [Fon56]. Australian [Bla56]. Automatic [Cab50, Rob52]. Autumn [Bax59, Cus55a, PS59]. Autumn-Spawned [Cus55a]. Auxiliary [Nie50]. Avoidance [Bal51].

B [Let52]. Bactericidal [VR54]. Bactériologie [Pic58]. Bacteriology [Pic58]. Balance [Nie58]. Baltic [Alm58, Chr50, DM58, Dem57, Dol56, Mol55, Pal51, Sim51a, Sjö58]. Baltique [Dol56]. Bank [Cha55]. Bann [Fro50, Low51]. Barents [Cor58, Tro58]. Basis [Bir58]. Bay [Let51]. Bays [Gib59a]. Beds [Han54]. Behavior [Fon53]. Behaviour [Bal51, Dee58]. belly [Wul56]. Benthos [Lon59]. Bering [Mos54]. bermudensis [Men59]. between [Nie58]. Bevan [Luc57]. Bibliography [Ano50a, Ano50b, Ano51a, Ano51b, Ano52a, Ano52b, Ano53a, Ano54a, Ano54b, Ano54c, Ano55a, Ano56a, Ano56b, Ano56c, Ano56d, Ano57a, Ano57b, Ano57c, Ano58a, Ano58b, Ano58c, Ano59a, Ano59b, Fri59b, Mar53, Mar55a, Sim50a, Sim53a]. Biogeographical [Ras59]. Biological [Bar52, JN53]. Biomass [DM58]. Biometric [Néd54]. Biométrique [lG52b]. Biométriques [Néd54]. Biomettric [lG52b]. Biscay [Let51]. Bjørn [Mos58]. Black [Mau54, Kri59]. Blanc [Mau54]. Blegvad [Pou52]. blennoides [Har50b]. Bløden [Han55]. Blooming [Sve53]. bluefin [Néd54]. Board [Bie56]. Body [Gei54]. bogaraveo [Cou54]. Bones [FK59, Har50b, Men50]. Bottle [LB58, VR54]. Bottles [Cox54b, Pic58]. Bouteilles [Pic58]. Bridge [SB56]. Brood [Cha55, SG52]. Brood-Strength [Cha55]. Brünn [Har50b]. Brünnich [Cou54]. Buccinum [HU59]. Bumpus [PF57].

C [Mau52, Nie52, Nie59, RV54]. Côte [Let52, Cou54]. Côtes [Cou52a, For50, Mau52]. Cadenat [Mau54]. caerulea [Far58]. Calanus [WR57b, WR59]. Capacities [Par51]. Capelin [FT51b]. Capture [Tho53, Tho54, Tho59]. Caractéristiques [CF50]. Carapace [CM53]. Carbon [Nie52, eWR57]. Cardium [Col56a]. Cartographical [Jur50]. Catch [Tho58]. Catches [Alm58, Bri56, Bur58, Sjö58]. caught [Fra53]. Causes [Sjö58]. Cell [Rob50]. Central [Nie59]. Changes [Bal51, Bla56, WR57a]. Characteristics [CF50]. Charlottenlund [Nie59]. Charts [Lum56, Wyr56]. Cheek [Cab50]. Chemical [Dan56]. Chemistry [BR54]. Chlorine [Men54]. Chlorinité [Men54]. Chlorinity [BR53, BR54, Her51, Kin50]. cinerea [Han54]. Circulation [Bow58]. Clam [Tay59]. Clark [Rus51]. Clarke [PF57]. Clarke-Bumpus [PF57]. Class [Cor58]. Closely [Gib59a]. Closing [BEL59]. cloustoni [WR57a]. Clupea [FT51b]. Coalfish [Sun57]. Coast [Cou52a, FT51a, Let52, Mau52, Cou54, Jur50, Sal57, Sjö58]. Coastal [Grø58b]. Coasts [Seg51, WR57a]. Coccoliths [HM54]. Cockles [Col56a]. Cod [BT50, Chr50, Coh59, Cor58, Dan53, Dan56, Dem57, Hol59, Mar57, Sun57, Tay58, Tro58, Wis58]. Cold [Dav53, Sim53b]. Colds [AE58]. Colin [Luc57]. Collection [HW56]. Collector [Rob52]. Collectors [Wal56]. Colorimetric [Rob50]. Colour [Jer54]. Comments [WR59]. Commercial [Bai55, Col56a]. Commission [Ano57e]. Committee [Ano57e, HKK^{+}59]. Common [Coh59]. Comparative [CH54, Mau54, Mau54]. Comparator [Rob50]. Comparing [Bir58]. Comparison [Kin50, PJP56, Rae51, Ras59, RV54]. Comparisons [GBMG54]. Competitive [Wal56]. Complex [Far58]. Complexes [Ras59]. Comportement [Fon53]. Composition [Dan56, Tho58]. Concept [Tai51]. Condition [Bai58, BGM59, Tho52, Bla51]. Conditions [Dav53, Sve53]. Conductivity [SB56]. Conseil [Søm51]. Considerations [Mor55, Woo56]. Considering [HKK^{+}59]. Content [BT50, Sim51a]. Contribution [CB53, Pos54, CB53, Pos54]. Convention [Ano57e]. Cooling [Sim51a]. Copepodite [WR59]. Copepodites [WR57b]. Copper [Atk57]. Cornish [Tho54]. Corps [Gei54]. Correlations [Gul53a]. Council [Kor50, Ano52c]. Counter [War58a]. counts [Bos59]. Crabs [Tho53, Tho54, Tho59]. Creel [Tho54, Tho59]. Critical [Mar56]. Crop [Bra58]. Crustacea [Mau58]. Cube [AP58]. Culture [KJ51, San57]. Current [Ano50a, Ano50b, Ano51a, Ano51b, Ano52a, Ano52b, Ano53a, Ano54a, Ano54b, Ano54c, Ano55a, Ano56a, Ano56b, Ano56c, Ano56d, Ano57a, Ano57b, Ano57c, Ano58a, Ano58b, Ano58c, Ano59a, Ano59b, Fri59b, Jon57, Mar53, Mar55a, Sim50a, Sim53a]. Cutter [Per50]. Cuvier [For50]. Cycle [Lis57]. Cycles [Woo58].

D [Cad52]. d’Afrique [Let52]. Dahl [Bag58, Søm52]. Dakar [Cad52]. Dangers [Fon56]. dans [Fon53]. Dark [VR54]. Data [CB59, GH59, Hod58, Néd54]. Day [Bag58, Bri56]. Daylight [AP58, Jer54]. Decadal [Sch52]. Demonstrate [RG58]. Denitrification [Bra59]. Denmark [Nie59]. Densities [Lon59]. Density [Cox57]. Depressor [Col59]. Depth [Ham55, Whi57]. Description [RG58, War58a]. Destruction [Han54, Wal56]. Determination [BR53, CN55, Føy50, FK59, Gei54, Her51, Kin50, LT56, Men54, Nie59, eWR57, vL57, Gei54, Men54, HKK^{+}59]. Determinations [Whi57]. Determining [Cox54a, Men50, Mos54, YH57]. deux [For50]. Device [CM53, War58b]. Devonshire [JS55]. Diatoms [Nic58]. Diet [Far58]. Diet-Induced [Far58]. Different [Rol53]. Diffusion [Ste57a]. Diphenylbenzidine [Atk54]. Direct [Atk57, SA59].

Discontinuity [Han51]. Discussion [Bag58]. Diseases [Opp53]. Dissolved [TK56]. Distribution [AP58, Qas56, Sim53b, Wal54]. Distributions [BW58]. Diurnal [Han51]. Dodd [Ano57f]. Donald [Car56]. Données [Néd54]. Dredge [Bai55]. Drift [vB55]. Drift-Nets [vB55]. d’un [Fra53]. d’une [Gei54, lG51]. Duration [Qas56, WR59]. during [And50, Bla56, Dav53, Ein52, FT51a, Sim53b]. Dutch [Dee54]. Dynamics [HGTK59, RG58, Cus59b].

E. [Gra54a]. Early [Mar56]. East [Bur58, CB59, Hod58]. Eastern [Bra58, Bra59]. Eau [Men54, Pic58]. Echo [Bal51, Cus52, Cus55b, WP50]. Echo-Sounding [Bal51, Cus55b, WP50]. Echo-Surveys [Cus52]. Economic [Jac54]. Eddy [Ste57a]. Edinburgh [Kor50]. Editor [Bag59, Bur50, Eal54, Low52, Ril53, Søm51, Wyr58]. Editorial [Ano50c, Ano50d, Ano55b, Ano56d, Fri59b]. EDM. [WR57a]. edule [Col56a]. edulis [CB53]. Edw [Tho55]. Edward [Wim58]. Eel [Dee54, Fro50]. Eels [Dee57, Fro50]. Effect [Cus59a, DO56, Kip57]. Effects [VR54]. Efficiency [Bri58, Tho53, Tho54, vB55]. Eggs [FT51b]. Eingeweidefett [Wul56]. Ekman [Kul54]. Electric [Rob50]. Electrical [Cox55]. Electron [HM54]. Elver [Dee52]. Elvers [Dee58, Low51]. Emploi [Pic58, lG51]. Employé [BR58]. Employed [Tho53]. encrasicholus [CF50]. Énergie [Fon56]. Energy [Fon56]. Engineering [Jac54]. English [Tro58]. Engraulis [CF50]. Enumeration [HW56]. Environmental [Bal51]. Errata [Ano53b]. Erratum [Ano56e]. erythrinus [Cou54]. Escallop [Bai55, Gib59a]. Escallops [Gib53]. Esox [FK59]. Espèces [lG52a]. Essay [Kes50]. Essex [Han54]. Established [Ano57e]. Estimation [Aas58, Atk54, Atk57, GH59, Kip57, SA59]. État [Fra53]. Étude [lG52b, BR58, CB53, Gei54, Mau54, Pos54]. Europe [Fri59c]. European [Bla51]. Examination [Bie56]. Expérience [Let51]. Experiment [Let51]. Experiments [Boe56, Cus55b, Dan53, Rey59, VR54, WP50]. Exploitation [Col56a]. Exploited [RG58]. Exploration [Kor50]. Exposure [BD57]. Extinction [Jer54].

F [Dol52, FT51b, lG52b]. Factors [Boe56, DD50, Dee54, Low51, Pal51, Whi57, Woo57]. Faroe [Rae51]. Farran [Wen50a]. Fat [BT50, Woo58, Wul56]. Fauna [Seg51, Ano54d]. Faune [Ano54d]. Fecundities [Bax59]. Females [WR57b]. Fertilization [Bra58]. Fifth [WR59]. Filters [Jer54]. Filtration [CH54]. Finland [Seg51]. finmarchicus [WR59]. Finnish [Jur50, Sjö58]. Fish [Ano53c, BGM59, BD59, Bri56, Cab50, CLV53, Cus52, Cus55b, DD50, Dra58, Fon53, Gei54, MM54, Mor55, Opp53, RG58, Sim53b, lC54]. Fish-Tags [MM54]. Fisheries [Ano57e, Bla56, Fro50, Gul53a, Wes53a]. Fishery [And50, Bev53, Bur53, HGTK59, Hun53, Ras59]. Fishes [Mar56, Men50, Mos54]. Fishing [Cus59a, Par51, Rol53, Tho53, WP50]. Flatfish [Par51]. Flotation [Bir57]. Flounder [Mol55]. Fluctuations [Alm58, BT50]. Fluorescent [Woo55]. Fonction [Fon53]. Food [Han55]. Fortunes [CLV53]. Fourth [Ano57e, Cou54]. Francaises [For50]. Francois [Tes50]. Free [Far58]. Freezing [Sim51a]. French [AE58, Ano54d, BR58, Cad52, CF50, Cou52b, Cou52a, Cou54, CB53, Dol52, Dol56, FV50, Fon53, Fon56, Fra53, Gei54, Hel52, Let51, Let52, Mau52, Mau54, Men54, Néd54, Pic58, Pos54, lG51, lG52a, lG52b, lG52c]. Fresh [Dee58]. Frigate [Str59]. Froids [AE58]. Function [BD57, Fon53]. Further [Rey59].

Gall [Des54]. Garstang [Har50a]. Gascogne [Let51]. Gascoyne [Dob50]. Gauge [YH57]. Gear [Rol53, WP50]. Genre [lG52a]. Genus [lG52a]. George [Wen50a]. Georges [Cha55]. German [Wul56]. Germo [Dol52]. Germon [Fra53]. Gill [Kip57]. Gill-Net [Kip57]. Girth [Mar54a, Mar57]. Glass [Pic58]. Gmelin [Dol52]. Gold [Sal57]. Golfe [Let51]. Grab [Bir57]. Grabs [Bir58]. Gran [Bra56a]. Grands [AE58]. Great [AE58]. Ground [Han55]. Grounds [FT51a, Par51]. group [Woo59]. Growth [Col56a, CW59, Dee57, FK59, GH59, Har50b, Hol59, Men50, Men59, Tay58, Tay59]. Growth-Rate [Col56a, Hol59, Men50]. Guide [Ano53c]. Gulf [Bri58]. Gunn [WR59].

Haaken [Bra56a]. Haddock [Cha55, Jon59, Mar54a]. Hake [Fri59c, Gul56, Hel52, Let52, Mau54, Cad52, lG52b, lG52c]. Hakes [Mau52]. Halibut [DB52, McI53, Tho52]. Handed [War58b]. Hansen [Mos58]. Harald [Fje58, Pou52]. Hareng [Dol56]. harengus [FT51b]. Harold [Tai58]. Harvesting [Jac54]. Hasberg [Bra56a]. Hauls [Bar51]. Hawaiian [Str59]. Heat [Sim51a]. Heldt [Des57]. helgolandicus [WR57b]. Helland [Mos58]. Helland-Hansen [Mos58]. Helmuth [Ore52]. Henry [Dob50]. Heri [Des57]. Herings [Wul56]. Hermaphrodite [FV50]. Herring [Aas58, And50, Bal51, Bax59, Bol58, Bur58, Cus55a, Cus59a, CB59, Dem57, Dol56, Dra58, Ein52, FT51a, FT51b, Han55, Hod51, Hod58, Jon57, MR58, PS59, Sjö58, Woo58, Woo59, Wul56]. High [Bri58, Col59, Her51]. High-Speed [Bri58, Col59]. History [Mar56]. Hoc [Ano57e]. Holacanthus [Men59]. Homarus [Tho55]. Horizontal [Bar51]. Huîtres [AE58]. Hugh [Sou50]. Hyaline [Bra56b]. Hydrographic [Lum56]. Hydrographical [Dav53]. Hydrography [Gul53a]. Hypophyso [Fon53].

Ice [Jur50, Pal51, Sch52]. Iceland [FT51a, Sch52]. Icelandic [Ein52, FT51b]. Ichthyological [Ano54d]. Ichthyologique [Ano54d]. Ichthyométrie [lG51]. ichthyometry [lG51]. III [Bri58]. IJsselmeer [Dee57]. Illumination [AP58]. Immature [Bra56b]. Immersion [YH57]. Improved [Cox57, Mau58, War58a]. Improvements [PF57]. Incoming [Sma59]. Increase [Seg51]. Index [Ano50e, Ano51c, Ano53d, Ano54e, Ano55c, Ano56f, Ano57d, Ano58d, Ano59c]. Indicator [Fro57]. Indigenous [Wis58]. Indirect [Sma59]. Indonesian [Wes53a, Wyr56]. Induced [Far58]. Influence [Seg51, Sim51a]. Influencing [Boe56, Low51]. Inland [Dee54]. Inorganic [SA59]. Inshore [JS55]. Instance [Str59]. Instrument [Mau58, MM54]. International [Kin50, Kor50, lG51, Ano57e]. internationale [lG51]. Interpretation [Bra56b, CB59, Hod58]. Intervals [GH59]. intestinalis [CB53]. Introductory [Cou52a]. Investigations [Bla56, Chr50, Mol55, Nik58, Woo59]. Ireland [Fro50, Gib53, Gib59a, Low51]. Irish [BR54, Gib59b]. Iron [SA59]. Island [DO56]. Islands [Str59]. Ister [Wul56, Wul56].

J [Cor53a]. J. [Ano57f, Dol52]. Jacob [Hav55]. Jean [Des54]. Jespersen [Rus52]. Johan [Hav55]. John [Car56]. Journal [Søm51]. Jurva [Gra54e].

Kite [Col59]. Kite-Otter [Col59]. Knudsen [Men54, Tho50, vL57]. Knut [Søm52].

L. [Bra56b, Col56a, CB53, FT51b, FK59, Gib53, Gib59a, HU59, Mau54]. Lacquer [Gei54]. Lambert [Bou58]. Lamellibranchs [CH54]. Laminaria [WR57a]. Laminariaceae [Wal54]. lancea [For50]. lanceolatus [For50]. Laque [Gei54]. Large [Dra58]. Larger [Wic58]. Larvae [Bri56, DD50, Mor55]. Larval [Str59]. Lateral [Ste57a]. Layer [Han51]. Leander [Gib59b]. Lemon [Rae51]. Length [CM53, Kip57, Mar54a, Mar57]. Length-Girth [Mar54a, Mar57]. Lesauvage [For50]. Letter [Bur50, Eal54, Low52, Ril53, Wyr58]. Letters [Ano57g, Bag59]. leur [Fon53, Pic58]. Liebert [Tes50]. Life [Mar56]. Light [Bal51, Dra58, VR54]. Limnologie [Pic58]. Limnology [Pic58]. Linn. [CF50]. Linné [Cou54]. Linneus [Cou54]. Lissner [Ore52]. List [Dol56]. Liste [Dol56]. Live [Sun57]. Lobster [Tho55, Tho58]. Lobsters [Tho53, Tho54, Tho59]. Location [MM54]. Lofoten [Dan53, Rol53]. Lohm [HM54]. Louis [Bou58]. Lowestoft [Par51]. lucius [FK59].

M. [Men54]. Mackerel [JS55, Str59]. Macrovegetation [Grø57b, Grø58b]. Made [Bri58]. Magnetic [CN55]. Males [WR57b]. Mallotus [FT51b]. Management [Bur53, Hun53]. Manipulation [Bie56]. Marine [Bar52, Bla56, Bur53, HW56, Jac54, LT56, Mar56, Mor55, Opp53, Qas56, RV54, Tai51, Woo55, eWR57]. marinus [Bra56b]. Market [McI53]. Marking [Cab50, HU59]. Marks [Ano53c]. Maroc [CF50, Cou52a, Let52, Mau52]. Marocaine [Cou54]. Marquage [Let51]. Martin [Tho50]. Mass [DO56, Str59]. Matiére [BR58]. Matter [BR58]. Matthews [Car56]. Maturité [FV50, Fra53]. Maturity [FV50, Fra53]. Maurice [Dob50]. maximus [Gib53, Gib59a]. Means [Cox55, Nie50]. Measurement [Bai58, CM53, Cox55, SB56, Sma59, War58b]. Measurements [AP58, Nie59]. Measuring [Mau58, Nie52, PJP56, RV54]. Meat [BD57]. Mechanical [Nic58]. Mechanisms [Bra58]. mediterranea [HM54]. Mediterranean [BR58]. Méditerranée [BR58]. Meeting [Ano57e, Kor50]. Mer [Fra53, Men54, Néd54]. Mercury [YH57]. Merlu [Hel52, Let52, Mau54, lG52b, lG52c]. Merluccius [Mau54, lG52a]. Merlus [Cad52, Mau52]. Mers [Hel52]. Mesh [Boe56, Gra54b, Gul56, LRPP54, Mar54a, Mar55b, Mar57, PJP56, Sav58]. Mesh-Measuring [PJP56]. Meteorological [Pal51, Tai51]. Meter [Cox57]. Method [Bie56, Bur58, Jon59, Jur50, KJ51, Nie50, YH57, vL57]. Methods [HU59, HKK^{+}59, PJP56, RV54, Sma59, Tho53]. Microbiology [Woo55]. Microorganisms [Kri59]. Microscope [HM54]. Microscopic [Bie56]. Microscopy [Woo55]. Migrating [Dee58]. Migration [Dee52, Dee54, Han51]. Mill [Sou50]. Model [RG58]. Modification [Bag58, vL57]. Modified [Bri58]. molischi [HM54]. Molluscan [Woo57]. Molluscs [Kor50]. Monthly [Wyr56]. Morocco [CF50, Cou52a, Cou54, Let52, Mau52]. Morphological [Gei54, CF50]. Morphologie [Fon53]. Morphologique [Gei54]. Morphologiques [CF50]. Morphology [Fon53]. Mortality [Sim53b, Str59, Tho55]. Movements [JS55, Jon57]. Mr. [Men54]. Muds [eWR57]. Müll. [FT51b]. Multiplane [Col59]. Multiple [BEL59]. Mussel [Rey59]. Mussels [Bai58]. Mytilicola [CB53]. Mytilus [BD57, CB53].

Name [Coh59]. Nanoplankton [KJ51, YR59]. Nansen [LB58]. Nares [Ano57f]. Natural [Str59, TK56]. Net [Bri58, Cor53a, Kip57, YR59]. Nets [Sav58, vB55]. Neutral [CH54]. News [Ano57g]. Nickel [LT56]. Night [Bag58, Bri56]. Nitrate [Atk54, Atk57, SA59]. Nitrification [Bra59]. Noir [Mau54]. Nord [Ano54d, Dol56, Néd54]. North [Bra59, Cus59a, Dav53, PS59, Sim53b, Ste57a, And50, Ano54d, Bal51, Cus55a, Dol56, Han55, Néd54, San57, Sim51a, Woo58]. North-Western [San57]. Northern [Fro50, Low51, And50, PS59, Ste57a]. Norway [Dra58]. Norwegian [Aas58, Cor58, Sma59]. notamment [Gei54]. Notation [HGTK59]. Note [Ano50c, Ano50d, Atk54, Cou52b, Cou52a, Cou54, GBMG54, Har50b, Hel52, LB58, Lum56, McI53, Cad52, Cou52b, Let52]. Notes [Atk57, Gib59a, Kin50, Nik58, SA59]. Notice [Ano53e, Ano55b, Ano56d, Fri59b]. Nouveaux [Pic58]. Nutrition [Mor55].

O. [FT51b]. Obituary [Ano53e]. Observations [Bal51, Bev53, BD59, For50, Fro50, Gib59b, HM54, Sim53b, Tho55]. Occurrence [Gib59b, LT56]. Ocean [Cor53a]. Oceanic [Bow58, Cus59b]. Océanographie [Pic58]. Oceanography [Pic58]. Oceans [Fon56, Lis57, Nie54, Fon56, Ras59]. October [Kor50]. off [FT51a, Jur50, Lon59, Sal57, Sch52, Seg51]. Offprints [Ano52c]. Opaque [Bra56b]. Opening [BEL59]. Opening-and-Closing [BEL59]. Opercular [FK59]. Organic [BR58, Nie52, Nie54, eWR57]. Organique [BR58]. Organisms [LT56, Wal56]. Oro [Let52]. other [Men50, Sch52]. Otolith [PS59, Per50]. Otoliths [Bra56b, Dan56, Men50, Mos54]. Otter [Bag58, Col59]. Oxygen [RV54, TK56, TG57]. Oyster [Han54, Wal56].

Oyster-Spat [Wal56]. Oysters [AE58, Bai58, CW59, San57].

Pacific [Bra58, Bra59, DB52, Ras59, Tay59, Tho52]. Pagellus [Cou52b, Cou54]. Papers [Kor50]. Parameters [GH59]. parasité [CB53]. Parasites [Dol52, Dol56, Dol52]. parasitized [CB53]. Parr [FV50]. Part [BR54]. particular [Gei54]. parvenu [FV50]. Pecten [Gib53, Gib59a]. Pennant [Gib59b]. Perennial [WR57a]. Period [Mar56]. Permanent [Ano57e]. Pettersson [LB58]. Philip [Wen50a]. Phosphorus [SA59]. Photo [Rob50]. Photo-Electric [Rob50]. Photometer [AP58]. Photosynthesis [RV54, VR54]. Phycis [Har50b]. Phytoplankton [Cor53a, HW56, Nie58, Sve53, YR59]. Picking [Nic58]. Pike [FK59]. Pilchard [Bla51]. pilchardus [Bla51]. Pipette [Nic58]. Pituitary [Fon53]. Pituitary-Thyroidian [Fon53]. Plaice [Mol55]. Plankton [Bar51, BEL59, Bie56, Col59, Cor58, Cus53, Fro57, Jac54, Nie50, Nor50, PF57, Rob52, Sav58, Sim53b, YH57]. Pliers [Cab50]. Poissons [Fon53, Gei54]. Pollution [Woo57]. Polythene [Cox54b]. Poor [Sjö58]. population [Cus59b]. Populations [Bla56, Cus53]. Portable [MM54]. Post [Rae51]. Post-War [Rae51]. Pot [Tho54]. Potential [WR57b]. Potentiometric [BR53, Her51]. Poul [Rus52]. pour [BR58, Gei54, Men54]. Practice [Bur53]. Prawns [CM53]. Pre [Rae51]. Pre-War [Rae51]. Precipitate [Nor50]. Precise [BR53]. Préface [lG52c, lG52c]. Préliminaire [Hel52, Cou52a]. Preliminary [Bai55, Col56a, Dan53, KJ51, Hel52]. Presented [Kor50]. Primary [Kri59, Mar58, Nie59]. Principles [Bev53]. pris [Fra53]. Problem [CW59, Cus59b, lG52c]. Probleme [lG52c]. Problems [DB52]. Production [BR58, Kri59, Mar58, Nie52, Nie54, Nie59, Cus59b]. Productivity [Woo56]. Publications [Ano52c]. Published [GBMG54]. pulchra [HM54]. Purification [Rey59, Woo57]. Purposes [JN53].

Quantitative [BEL59, KJ51]. Quatriéme [Cou54]. Quick [CM53].

Races [Cus55a]. Radiation [Sma59]. Radio [Nie52]. Radio-active [Nie52]. Rafinesque [lG52a]. Rapid [War58b]. Rapports [Fon53]. Rate [BGM59, Col56a, Hol59, Men50]. Rates [CH54, Tho55]. Ratio [BD57, BR54, Tho55, WR59]. Razor [Tay59]. reached [FV50]. Reactions [Dra58]. Reactive [SA59]. Real [Bla56]. Recorder [Ham55]. Red [CH54]. Redfish [Bra56b]. Rees [Luc57]. Reference [Dra58]. Regarding [Mor55]. Région [Cad52, Cad52]. Regulation [Bev53, DB52]. Relation [Cha55, Col56a, Fon53, Han51, Jon57, Qas56, VR54]. Relationships [Kip57, Mar54a, Mar57]. Relative [YR59]. Remarks [FT51b]. Report [Ano57e, Bai55, HKK^{+}59]. Research [Hun53, Tai51]. Researches [Dem57]. Résultants [Fon56]. Resulting [Fon56]. Results [Dan53, Tro58]. Return [BGM59]. Returns [Jon59]. Review [Kes50, Kor50]. Reviews [Aa.50, Aas54a, Aas54b, Aas57, B.50c, B.50d, B.50e, B.50b, B.50a, B.50f, B.50g, B.50h, B.50i, B.50j, B.53, B.55, B.57, Bar53, Bar54, Ber56, Ble52, Bow54, Bra53, Bra57, Bru50, Bru57, C.52b, C.52a, C.52c, C.53, C.54, C.56, Car53, Car57a, Car57b, Car58, Cat50, Cla53, Col54, Col55, Col56b, Col57, Coo50, CC53, Cor53b, D.53, D.54a, D.54b, D.55, D.56, D.57, Dah50a, Dah50b, Dah50c, Dar57, Daw53, Dea50, Dee53a, Dee53b, Def54, Des50a, Des50b, Des53a, Des53b, Dev50, Die52, Die53, Die56, Dub59, F.51, F.54, F.56a, F.56b, F.56c, Fag57, Fra57, Fra58, Fri55, Fur50, G.55, G.57a, G.57b, Gil56, Gor53a, Gor50a, Gor50b, Gor53b, Gor53c, Gra54c, Gra54d, Gro50, Gro55, Gro57a, Gro58a, Gul53b, H.51b, H.51a, H.53, H.55a, H.55b, Har57, Hav54, Hav56, Hes53, J.53, Jär50, Jen55a, Jen55b, Jen58, Jer53, Jes50, Jon56, Jon54a, Jon54b, Jos54, K.53b, K.53c, K.53d]. Reviews [K.53a, K.54, K.55, K.57a, K.57b, K.57c, Kal56, KJ56a, KJ56b, KJ57, Knu54, Kor53a, Kor53b, Kor53c, Kor54, Kor56a, Kor56b, Kor56c, Kor58a, Kor58b, Kra54, Kra58, L.50, L.51, L.54a, L.54b, L.55a, L.55b, Lel54, Let50a, Let50b, Lin53, Lis54, Lis55, Lis56a, Lis56b, Lod52, M.52, M.54, M.55, Mac52a, Mac52b, Mac59, Mar54b, Mar59, Mcl53, Mcl56, Mol57a, Mol57b, Mos50a, Mos50b, Nie56, Øst56, P.53a, P.53b, P.53c, P.54b, P.54a, P.55, P.56a, P.56b, P.57, Par52, Par56, Par59, Pen50, Pen54, Pic50, Poo54, Pou50, Pou51a, Pou51b, Pro50, R.50, R.52, R.53, R.54, R.55a, R.55b, R.56, Ree50, Ree53, Ric54, Roc56, Ros52, Ros57, Rou51, Ruu55, S.51, S.55, S.57a, S.57c, S.57b, S.58a, S.58b, S.58d, S.58c, Sæl56a, Sæl56b, Sæl56c, Sæt56, Sæt58, Sæt59, Seg56, Sim50b, Sim51b, Sim54, Sim55, Søm50, Søm58a, Søm58b, Ste54, Ste57b, Sto58, Svä52]. Reviews [Swa57, T.50, T.57, Tai50, Th.57, Tom54, Tre50, Tre54, Tuc51, Tuc58, Tul52, Urs57, Urs54, Vau52, Vau53, Vau54, Ver50, Vib55, W.51b, W.51a, W.53a, W.53c, W.53b, W.54a, W.54b, W.55, W.56, Wau57, Wee54, Wee58, Wen57a, Wen57b, WL58, Wes53b, Wes55, Wib53, Wib54a, Wib54b, Wib55, Wüs54, Wüs55b, Wüs55a, Wüs58]. Rings [Bla51, MR58]. Rio [Let52]. Risso [Cou52b]. Risto [Gra54e]. River [Fro50, Low51, Two57, Wen50b, Wen53]. Robert [Sou50, Rus51, Wim58]. Role [Kri59, Sch52, Ste57a]. Rotating [War58a]. Rouge [Néd54]. Roundfish [LRPP54]. Routine [BR53, Bos59]. Runs [Low51]. Russell [Gra54a].

S [Gra54a]. S. [HM54]. Saline [TG57]. Salinity [CN55, Cox54a, Cox55, Cox57, Ham55, HKK^{+}59, SB56, Seg51, Wyr56, vL57]. Salmon [Alm58, FV50, Nik58, Two57, Wen50b, Wen53]. Sampler [BEL59, Grø57b, JN53, PF57]. Samplers [Col59]. Samples [Bir57, Nor50, Rob52]. Sampling [Bur58, McI53]. Sardina [Bla51]. Sardine [Let51]. Sardinops [Far58]. Saumon [FV50]. Savage [Wim58]. Say [Han54]. Scales [Bla51, FK59, MR58]. Schill [HM54]. Schools [Bol58]. Scientific [Coh59, Kor50]. Scotland [Wal54, WR57a]. Scottish [Tho54]. Sea [Atk54, Atk57, BR53, CN55, Cox54a, Dee52, Dee58, Føy50, Hel52, Her51, Jer54, Kin50, Kor50, LT56, Men54, Nie52, Nie58, Nie59, Nor50, Rob50, SB56, Sim51a, SA59, And50, Bal51, BR54, Cor58, Cus55a, Cus59a, Dav53, Fra53, Han55, Kri59, Mos54, Néd54, PS59, Sim53b, Ste57a, Tro58, Woo58].

Sea-Water [Nor50]. Season [And50, Qas56]. Seasonal [Alm58, Sma59, Tho58, Cus59b]. Sebastes [Bra56b]. Sedimentation [Nie50]. Sediments [LT56]. Seine [LRPP54]. Seiner [GBMG54]. Seines [Gra54b]. Selbie [Rus51]. Selection [Boe56, Gra54b, Gul56, Kip57, LRPP54, Mar54a, Mar55b, Mar57, Sav58]. Selectivity [Rol53, Tho59]. Self [Woo57]. Self-Purification [Woo57]. Semi [Cab50]. Semi-Automatic [Cab50]. senegalensis [Mau54]. Separation [WR57b, WR59]. September [Ano57e]. serratus [Gib59b]. ses [Fon53]. Several [Mos54]. Sex [Tho55, WR59]. Sexes [WR59]. Sexual [FV50, Fra53]. Sexuelle [FV50, Fra53]. Shallow [Grø57b, Grø58b]. Shannon [Two57, Wen50b, Wen53]. Shape [Bol58]. Shell [BD57]. Shellfish [Kor50, War58b, Woo57]. Ship [Bie56]. Shoals [Jon57]. Shrimps [CM53]. Significance [BW58, YR59]. Silver [Dee54, Fro50]. Simple [Per50, Bos59]. Simplification [Men54, Men54]. Simulate [Jer54]. Singapore [Wic58]. Single [War58b]. Single-Handed [War58b]. Size [Bla56]. Skagerak [And50]. Slot [Nie59]. Small [Mau58]. Soles [Mar55b, Rae51]. Solubility [TG57]. Solution [CH54]. Some [Bev53, Boe56, Bow58, Cus55b, Dan53, FT51b, Gib59b, Kin50, Mor55, Sim53b, Tho58, TK56, Jon59, Qas56, Sch52, Wes53a]. Sorting [Bir57]. Sound [YR59]. Sounding [Bal51, Cus55b, WP50]. South [FT51a]. Southern [Cus59a, Dav53, DM58, Mol55, Sim53b]. Southernmost [Wis58]. Space [BW58]. Spain [San57]. Sparidae [Cou52a, Cou54]. Sparidés [Cou52a, Cou54]. Spat [Han54, Wal56]. Spawned [Cus55a]. Spawners [Bax59]. Spawning [FT51a, PS59, Qas56]. Special [Dra58, Kor50]. Species [lG52a]. Speed [Bri58, Col59]. Speeds [BD59]. Spring [Bax59, Dra58, FT51a]. Stage [WR57b, WR59]. Stage-V [WR57b]. Standard [HKK^{+}59, HGTK59, Kin50]. Standing [Bra58]. State [Fra53]. States [Fri59c]. Stations [Cor53a]. Statistical [Ano57g, BW58]. Statistics [Bar52]. Steuer [CB53]. Stirrer [CN55]. Stock [Aas58, And50, Tho58]. Stocks [DB52, Rae51, RG58, Tho52]. Straits [Wic58]. Strength [Aas58, Cha55, Cor58]. Strengths [SG52]. Studies [Bow58, Cus53, Jur50, KJ51, Nie50]. Study [BR58, CH54, Col56a, Gei54, Nor50, Opp53, CB53, Mau54, Pos54, lG52b]. Stunted [CW59]. Sub [HKK^{+}59]. Sub-Committee [HKK^{+}59]. Subcutaneous [MM54, lC54]. subject [Cad52]. Submarine [AP58]. Sugars [Atk57]. Suggested [Bal51]. Sujet [Cad52]. Sulphate [BR54]. Summer [Bax59, PS59, Sjö58]. Sunlight [VR54]. Surface [Gei54, Wyr56]. Survey [Sal57, Wic58]. Surveys [Cus52]. Survival [DD50]. Sverdrup [Fje58]. Swedish [Mol55]. Swimming [BD59]. Synoptic [Lum56]. Syracosphaera [HM54]. Systematics [KJ51].

Tåning [Fri59a]. Table [Men54]. Tag [BGM59, Cab50, lC54]. Tagged [BGM59, Jon59]. Tagging [BGM59, Dan53, Fri59c, Gib53, Let51, Tro58]. Tags [MM54]. Technique [Bir57, lG51]. Teleosts [Qas56]. Temperature [Ham55, Hol59, Tai51, Tay58, Tay59]. Temperature-Salinity-Depth [Ham55]. Temperatures [Cha55, JS55, Sch52]. Terminology [HGTK59]. Tesch [Hav55]. Tests [Bri58, BW58]. their [Fon53, Mar54a, Mar57, Pic58, Qas56, Ras59]. Theory [Bal51, Bur53]. Thermocline [Bra58]. Thermometric [Whi57]. Thompson [Tai58]. Thon [Néd54]. Thonidae [lG51]. Thonidés [Pos54, lG51]. Three [Gib59a]. Thyroidian [Fon53]. Thyroïdienne [Fon53]. Tidal [BD57, Jon57, Lis57]. Time [GH59, Qas56, Sim51a]. Titration [Men54]. Topography [Bra58]. Torbay [JS55]. Total [AP58]. Tow [Bri58]. Tow-Net [Bri58]. Towing [BEL59]. Transformations [Bar52]. Transmission [Cox54b]. Transport [Sun57]. Trawl [Bag58]. Trawler [GBMG54]. Trawler/ [GBMG54]. Trawling [Sal57]. Trawls [Boe56, Gra54b, Gul56, Mar55b, Par51]. Trials [Gra54b]. Tropical [Bra58, Bra59, Pos54, Lon59, Pos54]. tuna [Néd54, Pos54]. Tunisian [Hel52]. Tunisiennes [Hel52]. Turt [Dee52, Dee58]. Type [Bai55]. Types [PS59, Pic58]. Tyrrhenian [Fra53]. Tyrrhénienne [Fra53].

U.S.S.R. [Dem57]. Ulrik [Fje58]. Ultraplankton [KJ51]. undatum [HU59]. Underwater [Grø57b, Grø58b]. Unequal [GH59]. United [Fri59c]. upon [Seg51]. Urosalpinx [Han54]. Use [Atk54, Bar52, CN55, CH54, Fon56, LB58, Men50, Mos54, Nie52, Pic58, lG51, Gei54]. Used [BR58, Par51, RG58, Rol53]. USSR [Nik58]. Utilisation [Fon56, Gei54]. Utilization [Men59].

V [WR57b]. Vagn [Kul54]. Variability [Bag58]. Variation [Bri56, Far58, Cus59b]. Variations [Tho58]. Variétés [lG52a]. Varieties [lG52a]. Vedel [Fri59a]. Vernal [Sve53]. Verre [Pic58]. Vertical [BEL59]. Vigneron [Bag58]. Vigneron-Dahl [Bag58]. villosus [FT51b]. Vineyard [YR59]. Volume [Ano50e, Ano51c, Ano53d, Ano54e, Ano55c, Ano56f, Ano57d, Ano58d, Ano59c, Fro57, YH57]. vulgaris [Dee52, Dee58, Tho55].

Walbaum [Bla51]. Walfrid [Kul54]. Walter [Har50a]. War [Rae51]. Water [Atk54, Atk57, BR53, BT50, CN55, Cox54a, Cox54b, Dee58, Føy50, Her51, Hol59, JN53, Kin50, LT56, LB58, Men54, Nor50, Pic58, Rob52, SB56, Sim51a, SA59, TG57]. Water-Bottle [LB58]. Water-Sampler [JN53]. Waters [Dee54, FT51b, Gib59b, Grø57b, Grø58b, JS55, Sma59, TK56, Wyr56, YR59]. Weather [Cor53a]. Weight [Kip57]. West [Lon59, Gib59a]. Western [Cor58, San57]. Whelks [HU59]. White [Mau54]. Whiting [BGM59, Gul56, Mar54a, Mar57]. Willer [Büc54]. Winds [CLV53, Cha55, SG52]. Winter [Bax59, Dav53, Sim53b]. Work [Cox55, TK56]. World [Wis58].

XIX [Ano54e]. XVI [Ano50e]. XVII [Ano51c]. XVIII [Ano53d]. XX [Ano55c]. XXI [Ano56f]. XXII [Ano57d]. XXIII [Ano58d]. XXIV [Ano59c].

Year [Cor58, Ein52]. Year-Class [Cor58]. Years [Lis57].

Zinc [Atk57]. Zones [Bra56b, Dan56, Har50b]. Zoobenthic [DM58]. Zooplankton [Bra58, Han51, Nie58, War58a, Wic58, Bos59].

References

Aa:1950:R

[Aa.50] O. Aa. Reviews. ICES Journal of Marine Science, 17(1):68–69, October 1950. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/
17/1/68/627956.

Aasen:1954:Ra

[Aas54a] Olav Aasen. Reviews. ICES Journal of Marine Science, 19(3): 387–388, March 1954. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/19/3/387/662826.

Aasen:1954:Rb

[Aas54b] Olav Aasen. Reviews. ICES Journal of Marine Science, 19(3): 388–389, March 1954. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/19/3/388/662831.

Aasen:1957:R

[Aas57] Olav Aasen. Reviews. ICES Journal of Marine Science, 22(3): 345–346, June 1957. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/22/3/345/615296.

Aasen:1958:ESS

[Aas58] Olav Aasen. Estimation of the stock strength of the Norwegian herring. ICES Journal of Marine Science, 24(1):95–110, November 1958. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/
24/1/95/722084.

Aboul-Ela:1958:LGF

[AE58] Ibrahim Ahmed Aboul-Ela. L’action des grands froids sur les huîtres. (French) [The action of great colds on oysters]. ICES Journal of Marine Science, 23(3):443–452, July 1958. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL

http://academic.oup.com/icesjms/article/23/3/443/648839.

Alm:1958:SFC

[Alm58] Gunnar Alm. Seasonal fluctuations in the catches of salmon in the Baltic. ICES Journal of Marine Science, 23(3):399–433, July 1958. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/23/
3/399/648825.

Andersson:1950:HFS

[And50] K. A. Andersson. The herring fishery and the stock of herring in the northern North Sea and the Skagerak during the season 1949–50. ICES Journal of Marine Science, 17(1):40–43, October 1950. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/17/
1/40/627939.

Anonymous:1950:CBa

[Ano50a] Anonymous. Current bibliography. ICES Journal of Marine Science, 16(2):258–290, January 1950. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/16/2/258/668665.

Anonymous:1950:CBb

[Ano50b] Anonymous. Current bibliography. ICES Journal of Marine Science, 16(3):399–415, July 1950. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/16/3/399/607880.

Anonymous:1950:ENa

[Ano50c] Anonymous. Editorial note. ICES Journal of Marine Science, 16 (3):384, July 1950. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/16/3/384/607822.

Anonymous:1950:ENb

[Ano50d] Anonymous. Editorial note. ICES Journal of Marine Science, 17 (1):60, October 1950. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/17/1/60/627943.

Anonymous:1950:IVX

[Ano50e] Anonymous. Index to volume XVI. ICES Journal of Marine Science, 16(3):416–419, July 1950. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/16/3/416/607885.

Anonymous:1951:CBa

[Ano51a] Anonymous. Current bibliography. ICES Journal of Marine Science, 17(2):204–220, April 1951. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/17/2/204/555634.

Anonymous:1951:CBb

[Ano51b] Anonymous. Current bibliography. ICES Journal of Marine Science, 17(3):325–339, August 1951. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/17/3/325/628714.

Anonymous:1951:IVX

[Ano51c] Anonymous. Index of volume XVII. ICES Journal of Marine Science, 17(3):340–342, August 1951. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/17/3/340/628718.

Anonymous:1952:CBa

[Ano52a] Anonymous. Current bibliography. ICES Journal of Marine Science, 18(1):82–97, April 1952. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/18/1/82/641223.

Anonymous:1952:CBb

[Ano52b] Anonymous. Current bibliography. ICES Journal of Marine Science, 18(2):256–273, August 1952. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/18/2/256/887497.

Anonymous:1952:OCP

[Ano52c] Anonymous. Offprints from the Council’s publications. ICES Journal of Marine Science, 18(1):72, April 1952. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/18/1/72/641198.

Anonymous:1953:CB

[Ano53a] Anonymous. Current bibliography. ICES Journal of Marine Science, 18(3):374–390, January 1953. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/18/3/374/695571.

Anonymous:1953:E

[Ano53b] Anonymous. Errata. ICES Journal of Marine Science, 18(3):358, January 1953. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/
18/3/358/695461.

Anonymous:1953:GFM

[Ano53c] Anonymous. A guide to fish marks. ICES Journal of Marine Science, 19(2):241, October 1953. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/19/2/241/647158.

Anonymous:1953:IVX

[Ano53d] Anonymous. Index to volume XVIII. ICES Journal of Marine Science, 18(3):391–393, January 1953. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/18/3/391/695574.

Anonymous:1953:ON

[Ano53e] Anonymous. Obituary notice. ICES Journal of Marine Science, 19(1):90, May 1953. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/19/1/90/678910.

Anonymous:1954:CBa

[Ano54a] Anonymous. Current bibliography. ICES Journal of Marine Science, 19(3):394–414, March 1954. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/19/3/394/662859.

Anonymous:1954:CBb

[Ano54b] Anonymous. Current bibliography. ICES Journal of Marine Science, 20(1):106–132, July 1954. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/20/1/106/664121.

Anonymous:1954:CBc

[Ano54c] Anonymous. Current bibliography. ICES Journal of Marine Science, 20(2):218–248, December 1954. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/20/2/218/780612.

Anonymous:1954:FIL

[Ano54d] Anonymous. Faune ichthyologique de l’atlantique nord. (French) [Ichthyological fauna of the North Atlantic]. ICES Journal of Marine Science, 20(1):8, July 1954. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/20/1/8/664145.

Anonymous:1954:IVX

[Ano54e] Anonymous. Index to volume XIX. ICES Journal of Marine Science, 19(3):415–417, March 1954. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/19/3/415/662862.

Anonymous:1955:CB

[Ano55a] Anonymous. Current bibliography. ICES Journal of Marine Science, 21(1):86–115, October 1955. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/21/1/86/630470.

Anonymous:1955:EN

[Ano55b] Anonymous. Editorial notice. ICES Journal of Marine Science, 20(3):254, May 1955. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/20/3/254/697567.

Anonymous:1955:IVX

[Ano55c] Anonymous. Index to volume XX. ICES Journal of Marine Science, 20(3):362–364, May 1955. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/20/3/362/697692.

Anonymous:1956:CBa

[Ano56a] Anonymous. Current bibliography. ICES Journal of Marine Science, 21(2):204–238, April 1956. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/21/2/204/782928.

Anonymous:1956:CBb

[Ano56b] Anonymous. Current bibliography. ICES Journal of Marine Science, 21(3):337–371, June 1956. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/21/3/337/644058.

Anonymous:1956:CBc

[Ano56c] Anonymous. Current bibliography. ICES Journal of Marine Science, 22(1):107–141, October 1956. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/22/1/107/721887.

Anonymous:1956:ENC

[Ano56d] Anonymous. Editorial notice: The current bibliography. ICES Journal of Marine Science, 21(2):200–203, April 1956. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL

http://academic.oup.com/icesjms/article/21/2/200/782922.

Anonymous:1956:E

[Ano56e] Anonymous. Erratum. ICES Journal of Marine Science, 22(1):106, October 1956. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/
22/1/106/721884.

Anonymous:1956:IVX

[Ano56f] Anonymous. Index to volume XXI. ICES Journal of Marine Science, 21(3):372–374, June 1956. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/21/3/372/644060.

Anonymous:1957:CBa

[Ano57a] Anonymous. Current bibliography. ICES Journal of Marine Science, 22(2):229–262, February 1957. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/22/2/229/612688.

Anonymous:1957:CBb

[Ano57b] Anonymous. Current bibliography. ICES Journal of Marine Science, 22(3):353–398, June 1957. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/22/3/353/615327.

Anonymous:1957:CBc

[Ano57c] Anonymous. Current bibliography. ICES Journal of Marine Science, 23(1):108–144, October 1957. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/23/1/108/801747.

Anonymous:1957:IVX

[Ano57d] Anonymous. Index to volume XXII. ICES Journal of Marine Science, 22(3):399–402, June 1957. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/22/3/399/615329.

Anonymous:1957:IFC

[Ano57e] Anonymous. International Fisheries Convention 1946: Report of the Ad Hoc Committee Established at the Fourth Meeting of the Permanent Commission, September 1955. ICES Journal of Marine Science, 23(1):7–37, October 1957. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/23/1/7/801806.

Anonymous:1957:JDN

[Ano57f] Anonymous. J. Dodd Nares. ICES Journal of Marine Science, 22(3):269, June 1957. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/22/3/269/615235.

Anonymous:1957:SNL

[Ano57g] Anonymous. Statistical news letters. ICES Journal of Marine Science, 23(1):6, October 1957. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/23/1/6/801794.

Anonymous:1958:CBa

[Ano58a] Anonymous. Current bibliography. ICES Journal of Marine Science, 23(2):271–316, April 1958. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/23/2/271/688909.

Anonymous:1958:CBb

[Ano58b] Anonymous. Current bibliography. ICES Journal of Marine Science, 23(3):469–514, July 1958. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/23/3/469/648863.

Anonymous:1958:CBc

[Ano58c] Anonymous. Current bibliography. ICES Journal of Marine Science, 24(1):163–208, November 1958. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/24/1/163/722030.

Anonymous:1958:IVX

[Ano58d] Anonymous. Index to volume XXIII. ICES Journal of Marine Science, 23(3):515–517, July 1958. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/23/3/515/648869.

Anonymous:1959:CBa

[Ano59a] Anonymous. Current bibliography. ICES Journal of Marine Science, 24(2):379–424, March 1959. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/24/2/379/745069.

Anonymous:1959:CBb

[Ano59b] Anonymous. Current bibliography. ICES Journal of Marine Science, 24(3):504–548, June 1959. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/24/3/504/677846.

Anonymous:1959:IVX

[Ano59c] Anonymous. Index to volume XXIV. ICES Journal of Marine Science, 24(3):549–551, June 1959. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/24/3/549/677852.

Atkins:1958:CPM

[AP58] W. R. G. Atkins and H. H. Poole. Cube photometer measurements of the angular distribution of submarine daylight and the total submarine illumination. ICES Journal of Marine Science, 23 (3):327–336, July 1958. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/23/3/327/648777.

Atkins:1954:NUD

[Atk54] W. R. G. Atkins. Note on the use of diphenylbenzidine for the estimation of nitrate in sea water. ICES Journal of Marine Science, 20(2):153–155, December 1954. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/20/2/153/780484.

Atkins:1957:DEA

[Atk57] W. R. G. Atkins. The direct estimation of ammonia in sea water, with notes on nitrate, copper, zinc, and sugars. ICES Journal of Marine Science, 22(3):271–277, June 1957. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/22/3/271/615237.

B:1950:Re

[B.50a] F. B. Reviews. ICES Journal of Marine Science, 16(3):389–390, July 1950. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/
16/3/389/607833.

B:1950:Rd

[B.50b] M. B. Reviews. ICES Journal of Marine Science, 16(2):253–254, January 1950. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/
16/2/253/668653.

B:1950:Ra

[B.50c] N. F. B. Reviews. ICES Journal of Marine Science, 16(2):242–243, January 1950. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/
16/2/242/668539.

B:1950:Rb

[B.50d] T. B. Reviews. ICES Journal of Marine Science, 16(2):246–247, January 1950. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/
16/2/246/668565.

B:1950:Rc

[B.50e] T. B. Reviews. ICES Journal of Marine Science, 16(2):247–248, January 1950. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/
16/2/247/668578.

B:1950:Rf

[B.50f] T. B. Reviews. ICES Journal of Marine Science, 16(3):390, July 1950. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/16/
3/390/607837.

B:1950:Rg

[B.50g] T. B. Reviews. ICES Journal of Marine Science, 16(3):390–391, July 1950. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/
16/3/390/607841.

B:1950:Rh

[B.50h] T. B. Reviews. ICES Journal of Marine Science, 16(3):391, July 1950. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/16/
3/391/607844.

B:1950:Ri

[B.50i] T. B. Reviews. ICES Journal of Marine Science, 16(3):392–393, July 1950. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/
16/3/392/607847.

B:1950:Rj

[B.50j] T. B. Reviews. ICES Journal of Marine Science, 16(3):393, July 1950. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/16/
3/393/607849.

B:1953:R

[B.53] K. F. B. Reviews. ICES Journal of Marine Science, 18(3):359, January 1953. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/
18/3/359/695467.

B:1955:R

[B.55] T. B. Reviews. ICES Journal of Marine Science, 21(1):77–78, October 1955. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/
21/1/77/630439.

B:1957:R

[B.57] L. K. B. Reviews. ICES Journal of Marine Science, 22(3):344–345, June 1957. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/
22/3/344/615292.

Bagenal:1958:AVA

[Bag58] T. B. Bagenal. An analysis of the variability associated with the vigneron-dahl modification of the otter trawl by day and by night and a discussion of its action. ICES Journal of Marine Science, 24(1):62–79, November 1958. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/24/1/62/722073.

Bagenal:1959:LE

[Bag59] T. B. Bagenal. Letters to the Editor. ICES Journal of Marine Science, 24(3):499–501, June 1959. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/24/3/499/677842.

Baird:1955:PRN

[Bai55] R. H. Baird. A preliminary report on a new type of commercial escallop dredge. ICES Journal of Marine Science, 20(3):290–294, May 1955. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/
20/3/290/697583.

Baird:1958:MCM

[Bai58] R. H. Baird. Measurement of condition in mussels and oysters. ICES Journal of Marine Science, 23(2):249–257, April 1958. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL

http://academic.oup.com/icesjms/article/23/2/249/688844.

Balls:1951:ECH

[Bal51] Ronald Balls. Environmental changes in herring behaviour: a theory of light avoidance, as suggested by echo-sounding observations in the North Sea. ICES Journal of Marine Science, 17(3):274–298, August 1951. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/
17/3/274/628687.

Barnes:1951:HPH

[Bar51] H. Barnes. Horizontal plankton hauls. ICES Journal of Marine Science, 17(2):133–139, April 1951. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/17/2/133/555582.

Barnes:1952:UTM

[Bar52] H. Barnes. The use of transformations in marine biological statistics. ICES Journal of Marine Science, 18(1):61–71, April 1952. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL

http://academic.oup.com/icesjms/article/18/1/61/641194.

Barlow:1953:R

[Bar53] E. W. Barlow. Reviews. ICES Journal of Marine Science, 18(3): 360–361, January 1953. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/18/3/360/695471.

Barnes:1954:R

[Bar54] H. Barnes. Reviews. ICES Journal of Marine Science, 19(3):383, March 1954. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/
19/3/383/662812.

Baxter:1959:FWS

[Bax59] I. G. Baxter. Fecundities of winter–spring and summer–autumn herring spawners. ICES Journal of Marine Science, 25(1):73–80, October 1959. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/
25/1/73/620820.

Baird:1957:RSM

[BD57] R. H. Baird and R. E. Drinnan. The ratio of shell to meat in Mytilus as a function of tidal exposure to air. ICES Journal of Marine Science, 22(3):329–336, June 1957. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/22/3/329/615263.

Blaxter:1959:OSS

[BD59] J. H. S. Blaxter and W. Dickson. Observations on the swimming speeds of fish. ICES Journal of Marine Science, 24(3):472–479, June 1959. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/
24/3/472/677829.

Be:1959:QMO

[BEL59] Allan W. H. Bé, Maurice Ewing, and Laurence W. Linton. A quantitative multiple opening-and-closing plankton sampler for vertical towing. ICES Journal of Marine Science, 25(1):36–46, October 1959. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/25/
1/36/620811.

Bertelsen:1956:R

[Ber56] E. Bertelsen. Reviews. ICES Journal of Marine Science, 22(1): 98–99, October 1956. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/22/1/98/722011.

Beverton:1953:SOP

[Bev53] R. J. H. Beverton. Some observations on the principles of fishery regulation. ICES Journal of Marine Science, 19(1):56–68, May 1953. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/19/
1/56/678895.

Beverton:1959:WTH

[BGM59] R. J. H. Beverton, J. A. Gulland, and A. R. Margetts. Whiting tagging: How the tag return rate is affected by the condition of fish when tagged. ICES Journal of Marine Science, 25(1):53–57, October 1959. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/
25/1/53/620817.

Bieri:1956:MME

[Bie56] Robert Bieri. A method for the microscopic examination and manipulation of plankton on board ship. ICES Journal of Marine Science, 22(1):38–41, October 1956. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/22/1/38/721909.

Birkett:1957:FTS

[Bir57] Leon Birkett. Flotation technique for sorting grab samples. ICES Journal of Marine Science, 22(3):289–292, June 1957. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL

http://academic.oup.com/icesjms/article/22/3/289/615247.

Birkett:1958:BCG

[Bir58] L. Birkett. A basis for comparing grabs. ICES Journal of Marine Science, 23(2):202–207, April 1958. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/23/2/202/688819.

Blackburn:1951:CRS

[Bla51] Maurice Blackburn. “Condition rings” on scales of the European pilchard, Sardina pilchardus (Walbaum). ICES Journal of Marine Science, 17(2):181–195, April 1951. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/17/2/181/555602.

Blackburn:1956:RAC

[Bla56] M. Blackburn. Real and apparent changes in size of marine animal populations during the Australian fisheries investigations. ICES Journal of Marine Science, 21(3):284–295, June 1956. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL

http://academic.oup.com/icesjms/article/21/3/284/644024.

Blegvad:1952:R

[Ble52] H. Blegvad. Reviews. ICES Journal of Marine Science, 18(1):74–75, April 1952. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/
18/1/74/641200.

Boerema:1956:SEF

[Boe56] L. K. Boerema. Some experiments on factors influencing mesh selection in trawls. ICES Journal of Marine Science, 21(2):175–191, April 1956. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/
21/2/175/782877.

Bolster:1958:SHS

[Bol58] G. C. Bolster. On the shape of herring schools. ICES Journal of Marine Science, 23(2):228–234, April 1958. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/23/2/228/688830.

Bossanyi:1959:SAR

[Bos59] J. Bossanyi. A simple apparatus for routine zooplankton counts. ICES Journal of Marine Science, 24(3):452–454, June 1959. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/24/
3/452/677823.

Boury:1958:LL

[Bou58] Maurice Boury. Louis Lambert 1886–1957. ICES Journal of Marine Science, 23(2):155–156, April 1958. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/23/2/155/688782.

Bowden:1954:R

[Bow54] K. F. Bowden. Reviews. ICES Journal of Marine Science, 20 (2):206–209, December 1954. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/20/2/206/780541.

Bowden:1958:SRS

[Bow58] K. F. Bowden. Some recent studies of oceanic circulation. ICES Journal of Marine Science, 23(3):453–461, July 1958. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL

http://academic.oup.com/icesjms/article/23/3/453/648842.

Bather:1953:PRP

[BR53] J. M. Bather and J. P. Riley. The precise and routine potentiometric determination of the chlorinity of sea water. ICES Journal of Marine Science, 18(3):277–286, January 1953. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/18/3/277/695413.

Bather:1954:CIS

[BR54] J. M. Bather and J. P. Riley. The chemistry of the Irish Sea. Part I. The sulphate–chlorinity ratio. ICES Journal of Marine Science, 20(2):145–152, December 1954. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/20/2/145/780449.

Brouardel:1958:AEM

[BR58] Jean Brouardel and Emile Rinck. Appareillage employé en Méditerranée pour l’Étude de la production de matiére organique. (French) [Apparatus used in the Mediterranean for the study of the production of organic matter]. ICES Journal of Marine Science, 24(1):10–15, November 1958. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/24/1/10/721949.

Braarud:1953:R

[Bra53] Trygve Braarud. Reviews. ICES Journal of Marine Science, 19 (1):99–100, May 1953. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/19/1/99/678944.

Braarud:1956:HHG

[Bra56a] T. Braarud. Haaken Hasberg Gran 1870–1955. ICES Journal of Marine Science, 21(2):121–124, April 1956. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/21/2/121/782843.

Bratberg:1956:IOH

[Bra56b] Erling Bratberg. On the interpretation of the opaque and hyaline zones in the otoliths of immature redfish (Sebastes marinus L.). ICES Journal of Marine Science, 22(1):66–74, October 1956. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/22/
1/66/721926.

Braarud:1957:R

[Bra57] Trygve Braarud. Reviews. ICES Journal of Marine Science, 23(1): 102–103, October 1957. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/23/1/102/801696.

Brandhorst:1958:TTZ

[Bra58] Wilhelm Brandhorst. Thermocline topography, zooplankton standing crop, and mechanisms of fertilization in the Eastern tropical Pacific. ICES Journal of Marine Science, 24(1):16–31, November 1958. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/24/
1/16/722019.

Brandhorst:1959:NDE

[Bra59] Wilhelm Brandhorst. Nitrification and denitrification in the Eastern Tropical North Pacific. ICES Journal of Marine Science, 25(1): 3–20, October 1959. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/25/1/3/620809.

Bridger:1956:DNV

[Bri56] J. P. Bridger. On day and night variation in catches of fish larvae. ICES Journal of Marine Science, 22(1):42–57, October 1956. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/22/
1/42/721914.

Bridger:1958:ETM

[Bri58] J. P. Bridger. On efficiency tests made with a modified Gulf III highspeed tow-net. ICES Journal of Marine Science, 23(3):357–365, July 1958. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/
23/3/357/648788.

Bruun:1950:R

[Bru50] Anton Fr. Bruun. Reviews. ICES Journal of Marine Science, 16 (3):397–398, July 1950. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/16/3/397/607876.

Bruun:1957:R

[Bru57] Anton Fr. Bruun. Reviews. ICES Journal of Marine Science, 22 (3):351, June 1957. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/22/3/351/615315.

Bogucki:1950:FWF

[BT50] M. Bogucki and P. Trzesinski. Fluctuations in the water and fat content of the cod. ICES Journal of Marine Science, 16(2):208–210, January 1950. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/
16/2/208/668500.

Buckmann:1954:AW

[Büc54] A. Bückmann. Alfred Willer 1889–1952. ICES Journal of Marine Science, 19(3):293–296, March 1954. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/19/3/293/662734.

Burkenroad:1950:LE

[Bur50] Martin D. Burkenroad. Letter to the Editor. ICES Journal of Marine Science, 17(1):61, October 1950. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/17/1/61/627944.

Burkenroad:1953:TPM

[Bur53] Martin D. Burkenroad. Theory and practice of marine fishery management. ICES Journal of Marine Science, 18(3):300–310, January 1953. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/18/
3/300/695426.

Burd:1958:AMS

[Bur58] A. C. Burd. An analysis of the method of sampling the East Anglian herring catches. ICES Journal of Marine Science, 24(1):80– 94, November 1958. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/24/1/80/722079.

Buchanan-Wollaston:1958:STS

[BW58] H. J. Buchanan-Wollaston. Statistical tests for significance applicable to distributions in space. ICES Journal of Marine Science, 23 (2):161–172, April 1958. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/23/2/161/688799.

C:1952:Rb

[C.52a] D. H. C. Reviews. ICES Journal of Marine Science, 18(1):76–77, April 1952. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/
18/1/76/641205.

C:1952:Ra

[C.52b] H. C. Reviews. ICES Journal of Marine Science, 18(1):75–76, April 1952. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/
18/1/75/641202.

C:1952:Rc

[C.52c] H. A. C. Reviews. ICES Journal of Marine Science, 18(2):249–250, August 1952. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/
18/2/249/887485.

C:1953:R

[C.53] D. H. C. Reviews. ICES Journal of Marine Science, 19(2):214–215, October 1953. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/
19/2/214/647126.

C:1954:R

[C.54] D. H. C. Reviews. ICES Journal of Marine Science, 20(1):94–95, July 1954. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/
20/1/94/664158.

C:1956:R

[C.56] D. H. C. Reviews. ICES Journal of Marine Science, 22(1):94, October 1956. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/
22/1/94/721974.

Cable:1950:CTM

[Cab50] Louella E. Cable. A cheek tag for marking fish, with semi-automatic pliers for application of tag. ICES Journal of Marine Science, 16(2): 185–191, January 1950. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/16/2/185/668485.

Cadenat:1952:DNA

[Cad52] J. Cadenat. D. Note au sujet des merlus de la région de Dakar. (French) [D. Note on the subject of hake from the Dakar region]. ICES Journal of Marine Science, 18(2):230–233, August 1952. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/18/
2/230/887464.

Carruthers:1953:R

[Car53] J. N. Carruthers. Reviews. ICES Journal of Marine Science, 19(2): 211–213, October 1953. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/19/2/211/647120.

Carruthers:1956:DJM

[Car56] J. N. Carruthers. Donald John Matthews 1873–1956. ICES Journal of Marine Science, 22(1):3–8, October 1956. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/22/1/3/721901.

Carruthers:1957:Ra

[Car57a] J. N. Carruthers. Reviews. ICES Journal of Marine Science, 22 (3):340–341, June 1957. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/22/3/340/615278.

Carruthers:1957:Rb

[Car57b] J. N. Carruthers. Reviews. ICES Journal of Marine Science, 23 (1):98, October 1957. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/23/1/98/801841.

Carruthers:1958:R

[Car58] J. N. Carruthers. Reviews. ICES Journal of Marine Science, 23(2): 262–263, April 1958. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/23/2/262/688871.

Cattley:1950:R

[Cat50] J. G. Cattley. Reviews. ICES Journal of Marine Science, 16(3): 394–395, July 1950. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/16/3/394/607853.

Couteaux-Bargeton:1953:CEM

[CB53] M. Couteaux-Bargeton. Contribution à l’Étude de Mytilus edulis L. parasité par Mytilicola intestinalis Steuer. (French) [Contribution to the study of Mytilus edulis L. parasitized by Mytilicola intestinalis Steuer]. ICES Journal of Marine Science, 19(1):80–84, May 1953. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/
19/1/80/678904.

Cushing:1959:IEA

[CB59] D. H. Cushing and A. C. Burd. The interpretation of the East Anglian herring data, 1924–1955. ICES Journal of Marine Science, 24(2):277–282, March 1959. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/24/2/277/745015.

Cooper:1953:R

[CC53] L. H. N. Cooper and R. I. Currie. Reviews. ICES Journal of Marine Science, 19(2):209–211, October 1953. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/19/2/209/647117.

Coupe:1950:CMA

[CF50] René Coupe and Jean Furnestin. Les caractéristiques morphologiques des anchois (Engraulis encrasicholus Linn.) du Maroc. (French) [Morphological characteristics of anchovies (Engraulis encrasicholus Linn.) from Morocco]. ICES Journal of Marine Science, 16(2):182–184, January 1950. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/16/2/182/668481.

Cole:1954:UNR

[CH54] H. A. Cole and B. T. Hepper. The use of neutral red solution for the comparative study of filtration rates of lamellibranchs. ICES Journal of Marine Science, 20(2):197–203, December 1954. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL

http://academic.oup.com/icesjms/article/20/2/197/780523.

Chase:1955:WTR

[Cha55] Joseph Chase. Winds and temperatures in relation to the broodstrength of Georges Bank haddock. ICES Journal of Marine Science, 21(1):17–24, October 1955. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/21/1/17/630401.

Chrzan:1950:IBC

[Chr50] F. Chrzan. Investigations on the Baltic cod. ICES Journal of Marine Science, 16(2):192–207, January 1950. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/16/2/192/668493.

Clarke:1953:R

[Cla53] Robert Clarke. Reviews. ICES Journal of Marine Science, 19(1): 103–105, May 1953. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/19/1/103/678865.

Carruthers:1953:WFF

[CLV53] J. N. Carruthers, A. L. Lawford, and V. F. C. Veley. Winds and fish fortunes. ICES Journal of Marine Science, 18(3):354–358, January 1953. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/
18/3/354/695455.

Cole:1953:DQA

[CM53] H. A. Cole and M. N. Mistakidis. A device for the quick and accurate measurement of carapace length in prawns and shrimps. ICES Journal of Marine Science, 19(1):77–79, May 1953. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL

http://academic.oup.com/icesjms/article/19/1/77/678901.

Christie:1955:UMS

[CN55] A. A. Christie and O. F. Newman. The use of a magnetic stirrer in the determination of the salinity of sea water. ICES Journal of Marine Science, 21(1):3–5, October 1955. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/21/1/3/630406.

Cohen:1959:SNC

[Coh59] Daniel M. Cohen. The scientific name of the common cod. ICES Journal of Marine Science, 25(1):50–52, October 1959. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL

http://academic.oup.com/icesjms/article/25/1/50/620814.

Cole:1954:R

[Col54] H. A. Cole. Reviews. ICES Journal of Marine Science, 20(2): 217, December 1954. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/20/2/217/780606.

Cole:1955:R

[Col55] H. A. Cole. Reviews. ICES Journal of Marine Science, 21(1):85, October 1955. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/
21/1/85/630464.

Cole:1956:PSG

[Col56a] H. A. Cole. A preliminary study of growth-rate in cockles (Cardium edule L.) in relation to commercial exploitation. ICES Journal of Marine Science, 22(1):77–90, October 1956. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/22/1/77/721941.

Cole:1956:R

[Col56b] H. A. Cole. Reviews. ICES Journal of Marine Science, 21(3): 334–335, June 1956. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/21/3/334/644056.

Colman:1957:R

[Col57] John S. Colman. Reviews. ICES Journal of Marine Science, 22 (3):343–344, June 1957. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/22/3/343/615287.

Colton:1959:MKO

[Col59] John B. Colton, Jr. The multiplane kite-otter as a depressor for high-speed plankton samplers. ICES Journal of Marine Science, 25(1):29–35, October 1959. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/25/1/29/620808.

Cooper:1950:R

[Coo50] L. H. N. Cooper. Reviews. ICES Journal of Marine Science, 16 (3):385–387, July 1950. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/16/3/385/607824.

Corlett:1953:NPO

[Cor53a] John Corlett. Net phytoplankton at ocean Weather Stations “I” and “J”. ICES Journal of Marine Science, 19(2):178–190, October 1953. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/
19/2/178/647057.

Corlett:1953:R

[Cor53b] John Corlett. Reviews. ICES Journal of Marine Science, 19(2): 215–216, October 1953. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/19/2/215/647131.

Corlett:1958:PWB

[Cor58] John Corlett. Plankton in the Western Barents Sea and the yearclass strength of the arcto-Norwegian cod. ICES Journal of Marine Science, 23(3):354–356, July 1958. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/23/3/354/648786.

Coupe:1952:NPS

[Cou52a] René Coupé. Note préliminaire sur les Sparidés des côtes du Maroc. (French) [Introductory note on the Sparidae of the coast of Morocco]. ICES Journal of Marine Science, 18(1):37, April 1952. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/18/
1/37/641183.

Coupe:1952:NPA

[Cou52b] René Coupé. Note sur Pagellus acarne (Risso). (French) [Note on Pagellus acarne (Risso)]. ICES Journal of Marine Science, 18(1): 38–41, April 1952. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/18/1/38/641188.

Coupe:1954:QNS

[Cou54] R. Coupé. Quatriéme note sur les Sparidés de la Côte marocaine, Pagellus erythrinus Linné et Pagellus bogaraveo Brünnich. (French) [Fourth note on the Sparidae of the Morocco Coast, Pagellus erythrinus Linneus and Pagellus bogaraveo Brünnich]. ICES Journal of Marine Science, 19(3):363–367, March 1954. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL

http://academic.oup.com/icesjms/article/19/3/363/662751.

Cox:1954:ADS

[Cox54a] R. A. Cox. An apparatus for determining the salinity of sea water. ICES Journal of Marine Science, 20(1):9–17, July 1954. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL

http://academic.oup.com/icesjms/article/20/1/9/664152.

Cox:1954:WTP

[Cox54b] Roland A. Cox. Water transmission of polythene bottles. ICES Journal of Marine Science, 19(3):297–300, March 1954. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL

http://academic.oup.com/icesjms/article/19/3/297/662735.

Cox:1955:RWM

[Cox55] Roland A. Cox. Recent work on the measurement of salinity by electrical means. ICES Journal of Marine Science, 20(3):306–309, May 1955. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/
20/3/306/697590.

Cox:1957:ISD

[Cox57] Roland A. Cox. An improved salinity and density meter. ICES Journal of Marine Science, 23(1):38–46, October 1957. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL

http://academic.oup.com/icesjms/article/23/1/38/801766.

Cushing:1952:ESF

[Cus52] D. H. Cushing. Echo-surveys of fish. ICES Journal of Marine Science, 18(1):45–60, April 1952. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/18/1/45/641191.

Cushing:1953:SPP

[Cus53] D. H. Cushing. Studies on plankton populations. ICES Journal of Marine Science, 19(1):3–22, May 1953. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/19/1/3/678888.

Cushing:1955:ASH

[Cus55a] D. H. Cushing. On the autumn-spawned herring races of the North Sea. ICES Journal of Marine Science, 21(1):44–60, October 1955. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/
21/1/44/630408.

Cushing:1955:SES

[Cus55b] D. H. Cushing. Some echo-sounding experiments on fish. ICES Journal of Marine Science, 20(3):266–275, May 1955. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL

http://academic.oup.com/icesjms/article/20/3/266/697573.

Cushing:1959:EFH

[Cus59a] D. H. Cushing. On the effect of fishing on the herring of the Southern North Sea. ICES Journal of Marine Science, 24(2):283–307, March 1959. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/
24/2/283/745037.

Cushing:1959:SVO

[Cus59b] D. H. Cushing. The seasonal variation in oceanic production as a problem in population dynamics. ICES Journal of Marine Science, 24(3):455–464, June 1959. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/24/3/455/677826.

Cole:1959:PSG

[CW59] H. A. Cole and G. D. Waugh. The problem of stunted growth in oysters. ICES Journal of Marine Science, 24(2):355–365, March 1959. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/24/
2/355/745054.

D:1953:R

[D.53] C. L. D. Reviews. ICES Journal of Marine Science, 19(2):219, October 1953. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/
19/2/219/647142.

D:1954:Ra

[D.54a] C. L. D. Reviews. ICES Journal of Marine Science, 20(1):99, July 1954. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/20/
1/99/664171.

D:1954:Rb

[D.54b] C. L. D. Reviews. ICES Journal of Marine Science, 20(1):99–100, July 1954. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/
20/1/99/664173.

D:1955:R

[D.55] G. E. R. D. Reviews. ICES Journal of Marine Science, 21(1): 75–76, October 1955. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/21/1/75/630435.

D:1956:R

[D.56] J. D. Reviews. ICES Journal of Marine Science, 21(3):320–322, June 1956. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/
21/3/320/644033.

D:1957:R

[D.57] J. M. D. Reviews. ICES Journal of Marine Science, 22(2):226– 228, February 1957. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/22/2/226/612683.

Dahr:1950:Ra

[Dah50a] E. Dahr. Reviews. ICES Journal of Marine Science, 17(1):74, October 1950. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/
17/1/74/628018.

Dahr:1950:Rb

[Dah50b] E. Dahr. Reviews. ICES Journal of Marine Science, 17(1):74–75, October 1950. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/
17/1/74/628024.

Dahr:1950:Rc

[Dah50c] E. Dahr. Reviews. ICES Journal of Marine Science, 17(1):76, October 1950. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/
17/1/76/628030.

Dannevig:1953:TEC

[Dan53] Gunnar Dannevig. Tagging experiments on cod, Lofoten 1947–1952: Some preliminary results. ICES Journal of Marine Science, 19(2): 195–203, October 1953. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/19/2/195/647111.

Dannevig:1956:CCZ

[Dan56] Eva Henly Dannevig. Chemical composition of the zones in cod otoliths. ICES Journal of Marine Science, 21(2):156–159, April 1956. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/21/
2/156/782858.

Darbyshire:1957:R

[Dar57] J. Darbyshire. Reviews. ICES Journal of Marine Science, 23(1):97, October 1957. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/
23/1/97/801832.

David:1953:HCS

[Dav53] Vaux David. Hydrographical conditions in the Southern North Sea during the cold winter of 1946–1947. ICES Journal of Marine Science, 19(2):127–149, October 1953. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/19/2/127/647051.

Dawson:1953:R

[Daw53] E. Dawson. Reviews. ICES Journal of Marine Science, 19(1):91, May 1953. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/
19/1/91/678911.

Dunlop:1952:NPR

[DB52] H. A. Dunlop and F. H. Bell. New problems in regulation of Pacific halibut stocks. ICES Journal of Marine Science, 18(2):167–171, August 1952. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/
18/2/167/887439.

Dannevig:1950:FAS

[DD50] Alf Dannevig and Gunnar Dannevig. Factors affecting the survival of fish larvae. ICES Journal of Marine Science, 16(2):211–215, January 1950. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/
16/2/211/668505.

Deacon:1950:R

[Dea50] G. E. R. Deacon. Reviews. ICES Journal of Marine Science, 16(2): 237–239, January 1950. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/16/2/237/668523.

Deelder:1952:MEA

[Dee52] C. L. Deelder. On the migration of the elver (Anguilla vulgaris Turt.) at sea. ICES Journal of Marine Science, 18(2):187–218, August 1952. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/
18/2/187/887450.

Deelder:1953:Ra

[Dee53a] C. L. Deelder. Reviews. ICES Journal of Marine Science, 19(1):105, May 1953. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/
19/1/105/678871.

Deelder:1953:Rb

[Dee53b] C. L. Deelder. Reviews. ICES Journal of Marine Science, 19(1):106, May 1953. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/
19/1/106/678876.

Deelder:1954:FAM

[Dee54] C. L. Deelder. Factors affecting the migration of the silver eel in Dutch inland waters. ICES Journal of Marine Science, 20(2):177– 185, December 1954. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/20/2/177/780513.

Deelder:1957:GEI

[Dee57] C. L. Deelder. On the growth of eels in the IJsselmeer. ICES Journal of Marine Science, 23(1):83–88, October 1957. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL

http://academic.oup.com/icesjms/article/23/1/83/801819.

Deelder:1958:BEA

[Dee58] C. L. Deelder. On the behaviour of elvers (Anguilla vulgaris Turt.) migrating from the sea into fresh water. ICES Journal of Marine Science, 24(1):135–146, November 1958. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/24/1/135/721975.

Defant:1954:R

[Def54] A. Defant. Reviews. ICES Journal of Marine Science, 20(1):92–93, July 1954. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/
20/1/92/664154.

Dementjeva:1957:RUR

[Dem57] T. F. Dementjeva. Researches in the U.S.S.R. on Baltic herring and cod. ICES Journal of Marine Science, 22(3):309–321, June 1957. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/22/
3/309/615257.

Desbrosses:1950:Ra

[Des50a] P. Desbrosses. Reviews. ICES Journal of Marine Science, 16(2): 249–251, January 1950. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/16/2/249/668636.

Desbrosses:1950:Rb

[Des50b] P. Desbrosses. Reviews. ICES Journal of Marine Science, 16(2): 251–253, January 1950. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/16/2/251/668647.

Desbrosses:1953:Ra

[Des53a] P. Desbrosses. Reviews. ICES Journal of Marine Science, 18(3): 371–372, January 1953. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/18/3/371/695557.

Desbrosses:1953:Rb

[Des53b] P. Desbrosses. Reviews. ICES Journal of Marine Science, 18(3): 372–373, January 1953. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/18/3/372/695564.

Desbrosses:1954:JG

[Des54] P. Desbrosses. Jean le Gall 1894–1954. ICES Journal of Marine Science, 20(1):3–5, July 1954. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/20/1/3/664127.

Desbrosses:1957:HH

[Des57] P. Desbrosses. Heri Heldt (1891–1956). ICES Journal of Marine Science, 23(1):3–5, October 1957. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/23/1/3/801759.

Devold:1950:R

[Dev50] Finn Devold. Reviews. ICES Journal of Marine Science, 16(2):248, January 1950. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/
16/2/248/668590.

Dietrich:1952:R

[Die52] G. Dietrich. Reviews. ICES Journal of Marine Science, 18(2): 245–246, August 1952. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/18/2/245/887475.

Dietrich:1953:R

[Die53] G. Dietrich. Reviews. ICES Journal of Marine Science, 19(1):93–94, May 1953. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/
19/1/93/678927.

Dietrich:1956:R

[Die56] G. Dietrich. Reviews. ICES Journal of Marine Science, 21(3):319, June 1956. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/
21/3/319/644032.

Demel:1958:ZBS

[DM58] K. Demel and Z. Mulicki. The zoobenthic biomass in the Southern Baltic. ICES Journal of Marine Science, 24(1):43–54, November 1958. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/
24/1/43/722059.

Doty:1956:IME

[DO56] Maxwell S. Doty and Mikihiko Oguri. The Island mass effect. ICES Journal of Marine Science, 22(1):33–37, October 1956. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL

http://academic.oup.com/icesjms/article/22/1/33/721903.

Dobson:1950:HGM

[Dob50] A. T. A. Dobson. Henry Gascoyne Maurice. 1874–1950. ICES Journal of Marine Science, 17(1):3–6, October 1950. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL

http://academic.oup.com/icesjms/article/17/1/3/627931.

Dollfus:1952:PDG

[Dol52] Robert Ph. Dollfus. Parasites du Germo alalunga (J. F. Gmelin 1788). (French) [Parasites of Germo alalunga (J. F. Gmelin 1788)]. ICES Journal of Marine Science, 18(1):42–44, April 1952. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL

http://academic.oup.com/icesjms/article/18/1/42/641190.

Dollfus:1956:LPA

[Dol56] Robert Ph. Dollfus. Liste des parasites animaux du hareng de l’atlantique nord et de la baltique. (French) [List of animal parasites of North Atlantic and Baltic herring]. ICES Journal of Marine Science, 22(1):58–65, October 1956. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/22/1/58/721922.

Dragesund:1958:RFA

[Dra58] Olav Dragesund. Reactions of fish to artificial light, with special reference to large herring and spring herring in Norway. ICES Journal of Marine Science, 23(2):213–227, April 1958. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL

http://academic.oup.com/icesjms/article/23/2/213/688825.

Dubach:1959:R

[Dub59] Harold W. Dubach. Reviews. ICES Journal of Marine Science, 25(1):104–105, October 1959. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/25/1/104/620801.

Ealey:1954:LE

[Eal54] E. H. M. Ealey. Letter to the Editor. ICES Journal of Marine Science, 19(3):368, March 1954. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/19/3/368/662754.

Einarsson:1952:AIH

[Ein52] Hermann Einarsson. Analysis of the Icelandic herring during the year 1950. ICES Journal of Marine Science, 18(2):172–186, August 1952. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/
18/2/172/887443.

elWakeel:1957:DOC

[eWR57] S. K. el Wakeel and J. P. Riley. The determination of organic carbon in marine muds. ICES Journal of Marine Science, 22(2):180– 183, February 1957. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/22/2/180/612581.

F:1951:R

[F.51] J. H. F. Reviews. ICES Journal of Marine Science, 17(2):200–201, April 1951. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/
17/2/200/555623.

F:1954:R

[F.54] J. H. F. Reviews. ICES Journal of Marine Science, 19(3):377–378, March 1954. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/
19/3/377/662782.

F:1956:Ra

[F.56a] E. F. Reviews. ICES Journal of Marine Science, 21(3):323, June 1956. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/21/
3/323/644036.

F:1956:Rb

[F.56b] E. F. Reviews. ICES Journal of Marine Science, 21(3):323–324, June 1956. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/
21/3/323/644038.

F:1956:Rc

[F.56c] J. R. D. F. Reviews. ICES Journal of Marine Science, 22(1):91–92, October 1956. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/
22/1/91/721958.

Fage:1957:R

[Fag57] L. Fage. Reviews. ICES Journal of Marine Science, 23(1):100, October 1957. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/
23/1/100/801673.

Farris:1958:DIV

[Far58] David A. Farris. Diet-induced variation in the free amino acid complex of Sardinops caerulea. ICES Journal of Marine Science, 23(2): 235–244, April 1958. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/23/2/235/688835.

Fjeldstad:1958:HUS

[Fje58] J. E. Fjeldstad. Harald Ulrik Sverdrup 1888–1957. ICES Journal of Marine Science, 23(2):147–150, April 1958. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/23/2/147/688775.

Frost:1959:DAG

[FK59] Winifred E. Frost and Charlotte Kipling. The determination of the age and growth of pike (Esox lucius L.) from scales and opercular bones. ICES Journal of Marine Science, 24(2):314–341, March 1959. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/24/
2/314/745043.

Fontaine:1953:FHT

[Fon53] M. Fontaine. La fonction hypophyso–thyroïdienne des poissons dans ses rapports avec leur morphologie et leur comportement. (French) [The pituitary-thyroidian function of fish in its relation to their morphology and behavior]. ICES Journal of Marine Science, 19 (1):23–38, May 1953. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/19/1/23/678885.

Fontaine:1956:ODR

[Fon56] Maurice Fontaine. Les océans et les dangers résultants de l’utilisation de l’Énergie atomique. (French) [The oceans and dangers resulting from the use of atomic energy]. ICES Journal of Marine Science, 21(3):241–249, June 1956. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/21/3/241/644013.

Forest:1950:ODA

[For50] Jacques Forest. Observations sur deux ammodytidés des côtes françaises, Ammodytes lanceolatus Lesauvage et Ammodytes lancea Cuvier. ICES Journal of Marine Science, 16(2):179–181, January 1950. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/16/
2/179/668476.

Foyn:1950:ADS

[Føy50] Ernst Føyn. Ammonia determination in sea water. ICES Journal of Marine Science, 16(2):175–178, January 1950. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL

http://academic.oup.com/icesjms/article/16/2/175/668471.

Frade:1953:EMS

[Fra53] F. Frade. Sur l’État de maturité sexuelle d’un germon pris en mer tyrrhénienne. (French) [On the state of sexual maturity of an albacore caught in the Tyrrhenian Sea]. ICES Journal of Marine Science, 19(1):72–76, May 1953. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/19/1/72/678899.

Franc:1957:R

[Fra57] A. Franc. Reviews. ICES Journal of Marine Science, 23(1):100–101, October 1957. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/
23/1/100/801681.

Fraser:1958:R

[Fra58] J. H. Fraser. Reviews. ICES Journal of Marine Science, 24(1):160– 162, November 1958. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/24/1/160/722023.

Fridriksson:1955:R

[Fri55] Arni Fridriksson. Reviews. ICES Journal of Marine Science, 20 (3):323–324, May 1955. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/20/3/323/697652.

Fridriksson:1959:AVT

[Fri59a] Árni Fridriksson. Åge Vedel Tåning (1890–1958). ICES Journal of Marine Science, 24(2):211–214, March 1959. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/24/2/211/744982.

Fridriksson:1959:ENC

[Fri59b] Árni Fridriksson. Editorial notice: The current bibliography. ICES Journal of Marine Science, 24(3):497–498, June 1959. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL

http://academic.oup.com/icesjms/article/24/3/497/677840.

Fritz:1959:HTE

[Fri59c] Raymond L. Fritz. Hake tagging in Europe and the United States, 1931–1958. ICES Journal of Marine Science, 24(3):480–485, June 1959. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/24/
3/480/677830.

Frost:1950:EFR

[Fro50] Winifred E. Frost. The eel fisheries of the River Bann, Northern Ireland, and observations on the age of the silver eels. ICES Journal of Marine Science, 16(3):358–383, July 1950. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/16/3/358/607816.

Frolander:1957:PVI

[Fro57] Herbert F. Frolander. A plankton volume indicator. ICES Journal of Marine Science, 22(3):278–283, June 1957. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/22/3/278/615241.

Fridriksson:1951:HSG

[FT51a] A. Fridriksson and G. Timmermann. Herring spawning grounds off the South Coast of Iceland during spring 1950. ICES Journal of Marine Science, 17(2):172–180, April 1951. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/17/2/172/555598.

Fridriksson:1951:SRE

[FT51b] A. Fridriksson and G. Timmermann. Some remarks on the eggs of herring (Clupea harengus L.) and capelin (Mallotus villosus (O. F. Müll.)) in Icelandic waters. ICES Journal of Marine Science, 17(3):261–263, August 1951. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/17/3/261/628682.

Furnestin:1950:R

[Fur50] Jean Furnestin. Reviews. ICES Journal of Marine Science, 17(1): 69–71, October 1950. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/17/1/69/627961.

Fontaine:1950:PSH

[FV50] M. Fontaine and R. Vibert. Sur un parr de saumon hermaphrodite parvenu à maturité sexuelle. (French) [On a parr of hermaphrodite salmon that has reached sexual maturity]. ICES Journal of Marine Science, 16(2):216–226, January 1950. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/16/2/216/668511.

G:1955:R

[G.55] J. A. G. Reviews. ICES Journal of Marine Science, 20(3):330–332, May 1955. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/
20/3/330/697684.

G:1957:Ra

[G.57a] J. A. G. Reviews. ICES Journal of Marine Science, 22(2):221– 222, February 1957. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/22/2/221/612658.

G:1957:Rb

[G.57b] J. A. G. Reviews. ICES Journal of Marine Science, 22(2):222– 223, February 1957. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/22/2/222/612666.

Graham:1954:NPT

[GBMG54] M. Graham, R. J. H. Beverton, A. R. Margetts, and J. A. Gulland. A note on published Trawler/ seiner comparisons. ICES Journal of Marine Science, 20(1):51–55, July 1954. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/20/1/51/664130.

Geiger:1954:ULP

[Gei54] W. Geiger. Utilisation d’une laque pour l’Étude morphologique des poissons, notamment pour la détermination de la surface du corps. (French) [Use of a lacquer for the morphological study of fish, in particular for the determination of the surface of the body]. ICES Journal of Marine Science, 20(2):175–176, December 1954. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL

http://academic.oup.com/icesjms/article/20/2/175/780505.

Gulland:1959:EGP

[GH59] J. A. Gulland and S. J. Holt. Estimation of growth parameters for data at unequal time intervals. ICES Journal of Marine Science, 25(1):47–49, October 1959. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/25/1/47/620812.

Gibson:1953:TEP

[Gib53] F. A. Gibson. Tagging of escallops (Pecten maximus, L.) in Ireland. ICES Journal of Marine Science, 19(2):204–208, October 1953. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/19/
2/204/647115.

Gibson:1959:NEP

[Gib59a] F. A. Gibson. Notes on the escallop (Pecten maximus L.) in three closely associated bays in the West of Ireland. ICES Journal of Marine Science, 24(2):366–371, March 1959. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/24/2/366/745058.

Gibson:1959:SOO

[Gib59b] F. A. Gibson. Some observations on the occurrence of Leander serratus Pennant, in Irish waters. ICES Journal of Marine Science, 24(2):349–354, March 1959. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/24/2/349/745048.

Gilcrest:1956:R

[Gil56] Robert A. Gilcrest. Reviews. ICES Journal of Marine Science, 22(1):92–93, October 1956. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/22/1/92/721963.

Gordon:1950:Ra

[Gor50a] I. Gordon. Reviews. ICES Journal of Marine Science, 16(2):243– 244, January 1950. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/16/2/243/668546.

Gordon:1950:Rb

[Gor50b] I. Gordon. Reviews. ICES Journal of Marine Science, 16(2):244– 245, January 1950. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/16/2/244/668550.

Gorden:1953:R

[Gor53a] I. Gorden. Reviews. ICES Journal of Marine Science, 18(3):361– 362, January 1953. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/18/3/361/695476.

Gordon:1953:Ra

[Gor53b] I. Gordon. Reviews. ICES Journal of Marine Science, 18(3):362– 363, January 1953. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/18/3/362/695481.

Gordon:1953:Rb

[Gor53c] I. Gordon. Reviews. ICES Journal of Marine Science, 18(3):363– 364, January 1953. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/18/3/363/695490.

Graham:1954:R

[Gra54a] Michael Graham. E. S. Russell 1887–1954. ICES Journal of Marine Science, 20(2):135–139, December 1954. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/20/2/135/780423.

Graham:1954:TMS

[Gra54b] Michael Graham. Trials of mesh selection in trawls and seines. ICES Journal of Marine Science, 20(1):62–71, July 1954. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL

http://academic.oup.com/icesjms/article/20/1/62/664139.

Granqvist:1954:Ra

[Gra54c] Gunnar Granqvist. Reviews. ICES Journal of Marine Science, 20 (2):204–206, December 1954. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/20/2/204/780528.

Granqvist:1954:Rb

[Gra54d] Gunnar Granqvist. Reviews. ICES Journal of Marine Science, 20 (2):206, December 1954. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/20/2/206/780532.

Granqvist:1954:RJ

[Gra54e] Gunnar Granqvist. Risto Jurva 1888–1953. ICES Journal of Marine Science, 20(1):6–7, July 1954. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/20/1/6/664137.

Groen:1950:R

[Gro50] P. Groen. Reviews. ICES Journal of Marine Science, 16(2):239–241, January 1950. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/
16/2/239/668530.

Groen:1955:R

[Gro55] P. Groen. Reviews. ICES Journal of Marine Science, 20(3):313–315, May 1955. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/
20/3/313/697609.

Groen:1957:R

[Gro57a] P. Groen. Reviews. ICES Journal of Marine Science, 22(2): 214, February 1957. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/22/2/214/612624.

Grontved:1957:SUM

[Grø57b] Jul. Grøntved. A sampler for underwater macrovegetation in shallow waters. ICES Journal of Marine Science, 22(3):293–297, June 1957. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/22/
3/293/615251.

Groen:1958:R

[Gro58a] P. Groen. Reviews. ICES Journal of Marine Science, 23(2):266–267, April 1958. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/
23/2/266/688888.

Grontved:1958:UMS

[Grø58b] Jul. Grøntved. Underwater macrovegetation in shallow coastal waters. ICES Journal of Marine Science, 24(1):32–42, November 1958. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/24/
1/32/722052.

Gulland:1953:CFH

[Gul53a] J. A. Gulland. Correlations on fisheries hydrography. ICES Journal of Marine Science, 18(3):351–353, January 1953. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL

http://academic.oup.com/icesjms/article/18/3/351/695451.

Gulland:1953:R

[Gul53b] J. A. Gulland. Reviews. ICES Journal of Marine Science, 18(3): 365–366, January 1953. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/18/3/365/695502.

Gulland:1956:SHW

[Gul56] J. A. Gulland. On the selection of hake and whiting by the mesh of trawls. ICES Journal of Marine Science, 21(3):296–309, June 1956. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/21/
3/296/644028.

H:1951:Rb

[H.51a] S. J. H. Reviews. ICES Journal of Marine Science, 17(3):320–322, August 1951. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/
17/3/320/628709.

H:1951:Ra

[H.51b] T. J. H. Reviews. ICES Journal of Marine Science, 17(3):317–318, August 1951. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/
17/3/317/628693.

H:1953:R

[H.53] H. W. H. Reviews. ICES Journal of Marine Science, 19(1):94–95, May 1953. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/
19/1/94/678930.

H:1955:Ra

[H.55a] H. W. H. Reviews. ICES Journal of Marine Science, 20(3):310, May 1955. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/
20/3/310/697594.

H:1955:Rb

[H.55b] H. W. H. Reviews. ICES Journal of Marine Science, 20(3):310–311, May 1955. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/
20/3/310/697599.

Hamon:1955:TSD

[Ham55] B. V. Hamon. A temperature-salinity-depth recorder. ICES Journal of Marine Science, 21(1):72–73, October 1955. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/21/1/72/630425.

Hansen:1951:DMZ

[Han51] Kristian Vagn Hansen. On the diurnal migration of zooplankton in relation to the discontinuity layer. ICES Journal of Marine Science, 17(3):231–241, August 1951. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/17/3/231/628679.

Hancock:1954:DOS

[Han54] D. A. Hancock. The destruction of oyster spat by Urosalpinx cinerea (Say) on Essex oyster beds. ICES Journal of Marine Science, 20 (2):186–196, December 1954. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/20/2/186/780519.

Hansen:1955:FHB

[Han55] Vagn Hansen. The food of herring on the Bløden Ground (North Sea) in 1953. ICES Journal of Marine Science, 21(1):61–64, October 1955. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/
21/1/61/630418.

Hardy:1950:WG

[Har50a] A. C. Hardy. Walter Garstang. 1868–1949. ICES Journal of Marine Science, 17(1):7–12, October 1950. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/17/1/7/627963.

Hart:1950:NGZ

[Har50b] T. J. Hart. A note on growth zones in bones of Phycis blennoides Brünn. ICES Journal of Marine Science, 16(3):335–340, July 1950. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/16/
3/335/607807.

Hartley:1957:R

[Har57] W. G. Hartley. Reviews. ICES Journal of Marine Science, 22(3): 349–350, June 1957. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/22/3/349/615312.

Havinga:1954:R

[Hav54] B. Havinga. Reviews. ICES Journal of Marine Science, 20(1): 101–102, July 1954. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/20/1/101/664113.

Havinga:1955:JJT

[Hav55] B. Havinga. Johan Jacob Tesch. 1877–1954. ICES Journal of Marine Science, 20(3):251–253, May 1955. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/20/3/251/697565.

Havinga:1956:R

[Hav56] B. Havinga. Reviews. ICES Journal of Marine Science, 22(1): 105–106, October 1956. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/22/1/105/721875.

Heldt:1952:NPM

[Hel52] H. Heldt. E. note préliminaire sur le merlu des mers tunisiennes. (French) [E. Preliminary note on Tunisian sea hake]. ICES Journal of Marine Science, 18(2):234–235, August 1952. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/18/2/234/887465.

Hermann:1951:HAP

[Her51] Frede Hermann. High accuracy potentiometric determination of the chlorinity of sea water. ICES Journal of Marine Science, 17(3): 223–230, August 1951. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/17/3/223/628678.

Hessle:1953:R

[Hes53] Chr. Hessle. Reviews. ICES Journal of Marine Science, 18(3): 367–368, January 1953. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/18/3/367/695533.

Holt:1959:STN

[HGTK59] S. J. Holt, J. A. Gulland, C. Taylor, and S. Kurita. A standard terminology and notation for fishery dynamics. ICES Journal of Marine Science, 24(2):239–242, March 1959. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/24/2/239/744996.

Hermann:1959:RSC

[HKK^{+}59] F. Hermann, K. Kalle, F. Koczy, W. Maniece, and P. Tchernia. Report from the Sub-Committee for Considering Standard Methods for Salinity Determination. ICES Journal of Marine Science, 24(3): 429–445, June 1959. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/24/3/429/677819.

Halldal:1954:OCS

[HM54] Per Halldal and Joar Markali. Observations on coccoliths of Syracosphaera mediterranea Lohm., S. pulchra Lohm. and S. molischi Schill. in the electron microscope. ICES Journal of Marine Science, 19(3):329–336, March 1954. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/19/3/329/662741.

Hodgson:1951:HA

[Hod51] W. C. Hodgson. The herring atlas. ICES Journal of Marine Science, 17(2):196, April 1951. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/17/2/196/555608.

Hodgson:1958:IEA

[Hod58] W. C. Hodgson. On the interpretation of the East Anglian herring data 1923–1955. ICES Journal of Marine Science, 23(3):381–389, July 1958. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/
23/3/381/648815.

Holt:1959:WTC

[Hol59] S. J. Holt. Water temperature and cod growth-rate. ICES Journal of Marine Science, 24(2):374–376, March 1959. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/24/2/374/745065.

Hancock:1959:MMW

[HU59] D. A. Hancock and A. E. Urquhart. Methods for marking whelks (Buccinum undatum L.). ICES Journal of Marine Science, 24(3): 494–496, June 1959. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/24/3/494/677836.

Huntsman:1953:FMR

[Hun53] A. G. Huntsman. Fishery management and research. ICES Journal of Marine Science, 19(1):44–55, May 1953. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/19/1/44/678892.

Holmes:1956:ECM

[HW56] Robert W. Holmes and Theodore M. Widrig. The enumeration and collection of marine phytoplankton. ICES Journal of Marine Science, 22(1):21–32, October 1956. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/22/1/21/721894.

J:1953:R

[J.53] C. B. J. Reviews. ICES Journal of Marine Science, 18(3):369, January 1953. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/
18/3/369/695542.

Jackson:1954:EEA

[Jac54] Philip Jackson. Engineering and economic aspects of marine plankton harvesting. ICES Journal of Marine Science, 20(2):167–174, December 1954. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/
20/2/167/780499.

Jarvi:1950:R

[Jär50] T. H. Järvi. Reviews. ICES Journal of Marine Science, 17(1): 76–77, October 1950. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/17/1/76/628036.

Jensen:1955:Ra

[Jen55a] Aage J. C. Jensen. Reviews. ICES Journal of Marine Science, 21 (1):81, October 1955. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/21/1/81/630452.

Jensen:1955:Rb

[Jen55b] Aage J. C. Jensen. Reviews. ICES Journal of Marine Science, 21(1): 82–85, October 1955. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/21/1/82/630457.

Jensen:1958:R

[Jen58] Aa. J. C. Jensen. Reviews. ICES Journal of Marine Science, 23 (2):267–268, April 1958. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/23/2/267/688893.

Jerlov:1953:R

[Jer53] N. Jerlov. Reviews. ICES Journal of Marine Science, 19(2):213– 214, October 1953. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/19/2/213/647123.

Jerlov:1954:CFS

[Jer54] N. G. Jerlov. Colour filters to simulate the extinction of daylight in the sea. ICES Journal of Marine Science, 20(2):156–159, December 1954. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/
20/2/156/780490.

Jespersen:1950:R

[Jes50] Poul Jespersen. Reviews. ICES Journal of Marine Science, 16 (3):396, July 1950. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/16/3/396/607865.

Jensen:1953:WSB

[JN53] E. Aabye Jensen and E. Steemann Nielsen. A water-sampler for biological purposes. ICES Journal of Marine Science, 18(3):296– 299, January 1953. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/18/3/296/695422.

Jonsgaard:1954:Ra

[Jon54a] Åge Jonsgård. Reviews. ICES Journal of Marine Science, 20 (2):214–215, December 1954. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/20/2/214/780582.

Jonsgaard:1954:Rb

[Jon54b] Åge Jonsgård. Reviews. ICES Journal of Marine Science, 20 (2):215–216, December 1954. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/20/2/215/780591.

Jones:1956:R

[Jon56] R. Jones. Reviews. ICES Journal of Marine Science, 22(1):96, October 1956. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/
22/1/96/721992.

Jones:1957:MHS

[Jon57] F. R. Harden Jones. Movements of herring shoals in relation to the tidal current. ICES Journal of Marine Science, 22(3):322–328, June 1957. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/
22/3/322/615260.

Jones:1959:MAS

[Jon59] R. Jones. A method of analysis of some tagged haddock returns. ICES Journal of Marine Science, 25(1):58–72, October 1959. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/25/
1/58/620819.

Joseph:1954:R

[Jos54] Joachim Joseph. Reviews. ICES Journal of Marine Science, 19(3): 371–375, March 1954. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/19/3/371/662765.

Jackman:1955:TMM

[JS55] L. A. J. Jackman and G. A. Steven. Temperatures and mackerel movements in the inshore waters of torbay, devonshire. ICES Journal of Marine Science, 21(1):65–71, October 1955. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL

http://academic.oup.com/icesjms/article/21/1/65/630421.

Jurva:1950:CMS

[Jur50] Risto Jurva. On the cartographical method for studies of the ice off the Finnish Coast. ICES Journal of Marine Science, 16(3):293–306, July 1950. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/
16/3/293/607777.

K:1953:Rd

[K.53a] F. F. K. Reviews. ICES Journal of Marine Science, 19(1):97–99, May 1953. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/
19/1/97/678942.

K:1953:Ra

[K.53b] K. K. Reviews. ICES Journal of Marine Science, 19(1):92, May 1953. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/19/
1/92/678915.

K:1953:Rb

[K.53c] K. K. Reviews. ICES Journal of Marine Science, 19(1):92–93, May 1953. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/
19/1/92/678920.

K:1953:Rc

[K.53d] K. K. Reviews. ICES Journal of Marine Science, 19(1):93, May 1953. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/19/
1/93/678925.

K:1954:R

[K.54] K. K. Reviews. ICES Journal of Marine Science, 19(3):379, March 1954. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/19/
3/379/662790.

K:1955:R

[K.55] Ph. H. K. Reviews. ICES Journal of Marine Science, 21(1):74, October 1955. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/
21/1/74/630427.

K:1957:Ra

[K.57a] F. F. K. Reviews. ICES Journal of Marine Science, 22(2):209– 211, February 1957. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/22/2/209/612605.

K:1957:Rb

[K.57b] K. K. Reviews. ICES Journal of Marine Science, 22(3):337–338, June 1957. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/
22/3/337/615268.

K:1957:Rc

[K.57c] K. K. Reviews. ICES Journal of Marine Science, 22(3):338–339, June 1957. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/
22/3/338/615271.

Kalle:1956:R

[Kal56] K. Kalle. Reviews. ICES Journal of Marine Science, 21(3):324–325, June 1956. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/
21/3/324/644041.

Kesteven:1950:ER

[Kes50] G. L. Kesteven. Essay review. ICES Journal of Marine Science, 16(2):227–236, January 1950. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/16/2/227/668516.

King:1950:SND

[Kin50] J. King. Some notes on the determination of the chlorinity of sea water by comparison with the international standard. ICES Journal of Marine Science, 16(2):167–174, January 1950. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL

http://academic.oup.com/icesjms/article/16/2/167/668466.

Kipling:1957:EGN

[Kip57] C. Kipling. The effect of gill-net selection on the estimation of weight–length relationships. ICES Journal of Marine Science, 23 (1):51–63, October 1957. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/23/1/51/801786.

Knight-Jones:1951:PSN

[KJ51] E. W. Knight-Jones. Preliminary studies of nanoplankton and ultraplankton systematics and abundance by a quantitative culture method. ICES Journal of Marine Science, 17(2):140–155, April 1951. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/17/
2/140/555591.

Knight-Jones:1956:Ra

[KJ56a] E. W. Knight-Jones. Reviews. ICES Journal of Marine Science, 21 (3):333–334, June 1956. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/21/3/333/644053.

Knight-Jones:1956:Rb

[KJ56b] E. W. Knight-Jones. Reviews. ICES Journal of Marine Science, 21 (3):335–336, June 1956. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/21/3/335/644057.

Knight-Jones:1957:R

[KJ57] E. W. Knight-Jones. Reviews. ICES Journal of Marine Science, 22(2):219–220, February 1957. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/22/2/219/612652.

Knudsen:1954:R

[Knu54] J. Knudsen. Reviews. ICES Journal of Marine Science, 20(2):213– 214, December 1954. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/20/2/213/780574.

Korringa:1950:RPM

[Kor50] P. Korringa. A review of the papers on molluscs presented at the Special Scientific Meeting on Shellfish of the International Council for the Exploration of the Sea, Edinburgh, 10th October 1949. ICES Journal of Marine Science, 17(1):44–59, October 1950. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL

http://academic.oup.com/icesjms/article/17/1/44/627940.

Korringa:1953:Ra

[Kor53a] P. Korringa. Reviews. ICES Journal of Marine Science, 18(3): 369–371, January 1953. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/18/3/369/695550.

Korringa:1953:Rb

[Kor53b] P. Korringa. Reviews. ICES Journal of Marine Science, 19(2):220, October 1953. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/
19/2/220/647144.

Korringa:1953:Rc

[Kor53c] P. Korringa. Reviews. ICES Journal of Marine Science, 19(2): 220–222, October 1953. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/19/2/220/647147.

Korringa:1954:R

[Kor54] P. Korringa. Reviews. ICES Journal of Marine Science, 19(3): 392–393, March 1954. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/19/3/392/662854.

Korringa:1956:Ra

[Kor56a] P. Korringa. Reviews. ICES Journal of Marine Science, 21(2): 197–199, April 1956. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/21/2/197/782912.

Korringa:1956:Rb

[Kor56b] P. Korringa. Reviews. ICES Journal of Marine Science, 22(1): 101–102, October 1956. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/22/1/101/721855.

Korringa:1956:Rc

[Kor56c] P. Korringa. Reviews. ICES Journal of Marine Science, 22(1): 103–104, October 1956. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/22/1/103/721860.

Korringa:1958:Ra

[Kor58a] P. Korringa. Reviews. ICES Journal of Marine Science, 23(3): 466–467, July 1958. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/23/3/466/648852.

Korringa:1958:Rb

[Kor58b] P. Korringa. Reviews. ICES Journal of Marine Science, 23(3): 467–468, July 1958. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/23/3/467/648857.

Kramp:1954:R

[Kra54] P. L. Kramp. Reviews. ICES Journal of Marine Science, 20(2):212– 213, December 1954. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/20/2/212/780569.

Krauss:1958:R

[Kra58] W. Krauss. Reviews. ICES Journal of Marine Science, 24(1):156– 157, November 1958. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/24/1/156/721998.

Kriss:1959:RMP

[Kri59] A. E. Kriss. The role of microorganisms in the primary production of the Black Sea. ICES Journal of Marine Science, 24(2):221–230, March 1959. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/
24/2/221/744987.

Kullenberg:1954:VWE

[Kul54] B. Kullenberg. Vagn Walfrid Ekman. 1874–1954. ICES Journal of Marine Science, 20(2):140–144, December 1954. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL

http://academic.oup.com/icesjms/article/20/2/140/780425.

L:1950:R

[L.50] A. J. L. Reviews. ICES Journal of Marine Science, 16(2):241–242, January 1950. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/
16/2/241/668534.

L:1951:R

[L.51] K. L. Reviews. ICES Journal of Marine Science, 17(3):322–324, August 1951. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/
17/3/322/628710.

L:1954:Ra

[L.54a] C. E. L. Reviews. ICES Journal of Marine Science, 20(1):102–105, July 1954. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/
20/1/102/664118.

L:1954:Rb

[L.54b] E. L. Reviews. ICES Journal of Marine Science, 20(2):211–212, December 1954. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/
20/2/211/780562.

L:1955:Ra

[L.55a] E. L. Reviews. ICES Journal of Marine Science, 20(3):316, May 1955. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/20/
3/316/697622.

L:1955:Rb

[L.55b] E. L. Reviews. ICES Journal of Marine Science, 20(3):316, May 1955. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/20/
3/316/697627.

Lee:1958:NUN

[LB58] A. J. Lee and G. C. Baxter. A note on the use of the Nansen– Pettersson water-bottle. ICES Journal of Marine Science, 23(2): 157–160, April 1958. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/23/2/157/688785.

leCren:1954:STF

[lC54] E. D. le Cren. A subcutaneous tag for fish. ICES Journal of Marine Science, 20(1):72–82, July 1954. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/20/1/72/664141.

Leloup:1954:R

[Lel54] E. Leloup. Reviews. ICES Journal of Marine Science, 19(3): 379–380, March 1954. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/19/3/379/662796.

Letaconnoux:1950:Ra

[Let50a] R. Letaconnoux. Reviews. ICES Journal of Marine Science, 16 (2):245, January 1950. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/16/2/245/668555.

Letaconnoux:1950:Rb

[Let50b] R. Letaconnoux. Reviews. ICES Journal of Marine Science, 16(2): 245–246, January 1950. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/16/2/245/668560.

Letaconnoux:1951:EMS

[Let51] R. Letaconnoux. Une expérience de marquage sur la sardine du Golfe de Gascogne. (French) [A tagging experiment on the Bay of Biscay] sardine. ICES Journal of Marine Science, 17(3):264–266, August 1951. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/
17/3/264/628684.

Letaconnoux:1952:BNM

[Let52] R. Letaconnoux. B. Note sur le merlu de la côte d’Afrique. (Maroc et Rio de Oro). (French) [B. Note on the African coast hake. (Morocco and Rio de Oro)]. ICES Journal of Marine Science, 18(2):223, August 1952. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/
18/2/223/887460.

leGall:1951:ITE

[lG51] Jean le Gall. Ichthyométrie des Thonidés. De l’emploi d’une technique internationale. (French) [Thonidae ichthyometry. On the use of an international technique]. ICES Journal of Marine Science, 17(3):267–273, August 1951. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/17/3/267/628685.

leGall:1952:EVD

[lG52a] Jean le Gall. A. Les espèces et variétés du genre Merluccius Rafinesque. (French) [A. The species and varieties of the genus Merluccius Rafinesque]. ICES Journal of Marine Science, 18(2): 221–222, August 1952. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/18/2/221/887456.

leGall:1952:FEB

[lG52b] Jean le Gall. F. Étude biométrique du merlu. (French) [F. Biomettric study of the hake]. ICES Journal of Marine Science, 18(2): 236–240, August 1952. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/18/2/236/887468.

leGall:1952:PPD

[lG52c] Jean le Gall. Préface. Le “probleme du merlu”. (French) [Preface. The “hake problem”]. ICES Journal of Marine Science, 18(2): 219–220, August 1952. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/18/2/219/887454.

Lindstrom:1953:R

[Lin53] T. Lindström. Reviews. ICES Journal of Marine Science, 19(2): 218–219, October 1953. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/19/2/218/647138.

Lisitzin:1954:R

[Lis54] E. Lisitzin. Reviews. ICES Journal of Marine Science, 20(1):93–94, July 1954. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/
20/1/93/664155.

Lisitzin:1955:R

[Lis55] E. Lisitzin. Reviews. ICES Journal of Marine Science, 21(1):74–75, October 1955. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/
21/1/74/630430.

Lisitzin:1956:Ra

[Lis56a] E. Lisitzin. Reviews. ICES Journal of Marine Science, 21(3): 325–326, June 1956. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/21/3/325/644043.

Lisitzin:1956:Rb

[Lis56b] E. Lisitzin. Reviews. ICES Journal of Marine Science, 21(3): 326–327, June 1956. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/21/3/326/644045.

Lisitzin:1957:TCY

[Lis57] Eugenie Lisitzin. The tidal cycle of 18.6 years in the oceans. ICES Journal of Marine Science, 22(2):147–151, February 1957. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL

http://academic.oup.com/icesjms/article/22/2/147/612555.

Lodge:1952:R

[Lod52] Sheila Lodge. Reviews. ICES Journal of Marine Science, 18(2): 250–251, August 1952. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/18/2/250/887487.

Longhurst:1959:BDT

[Lon59] Alan R. Longhurst. Benthos densities off Tropical West Africa. ICES Journal of Marine Science, 25(1):21–28, October 1959. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/25/
1/21/620807.

Lowe:1951:FIR

[Low51] Rosemary H. Lowe. Factors influencing the runs of elvers in the River Bann, Northern Ireland. ICES Journal of Marine Science, 17(3):299–315, August 1951. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/17/3/299/628689.

Lowe:1952:LE

[Low52] Rosemary H. Lowe. Letter to the Editor. ICES Journal of Marine Science, 18(2):241–244, August 1952. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/18/2/241/887472.

Lucas:1954:MSR

[LRPP54] C. E. Lucas, A. Ritchie, B. B. Parrish, and J. A. Pope. Mesh selection in the roundfish seine. ICES Journal of Marine Science, 20(1):35–50, July 1954. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/20/1/35/664129.

Laevastu:1956:DON

[LT56] Taivo Laevastu and Thomas G. Thompson. The determination and occurrence of nickel in sea water, marine organisms, and sediments. ICES Journal of Marine Science, 21(2):125–143, April 1956. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/21/
2/125/782847.

Lucas:1957:CBR

[Luc57] C. E. Lucas. Colin Bevan Rees 1914–1956. ICES Journal of Marine Science, 22(3):267–269, June 1957. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/22/3/267/615232.

Lumby:1956:NSH

[Lum56] J. R. Lumby. A note on synoptic hydrographic charts. ICES Journal of Marine Science, 21(3):250–262, June 1956. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL

http://academic.oup.com/icesjms/article/21/3/250/644015.

M:1952:R

[M.52] N. A. M. Reviews. ICES Journal of Marine Science, 18(1):80–81, April 1952. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/
18/1/80/641220.

M:1954:R

[M.54] S. M. M. Reviews. ICES Journal of Marine Science, 19(3):378–379, March 1954. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/
19/3/378/662785.

M:1955:R

[M.55] A. R. M. Reviews. ICES Journal of Marine Science, 20(3):324–326, May 1955. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/
20/3/324/697656.

Macfarlane:1952:Ra

[Mac52a] P. R. C. Macfarlane. Reviews. ICES Journal of Marine Science, 18(2):253–254, August 1952. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/18/2/253/887491.

Macfarlane:1952:Rb

[Mac52b] P. R. C. Macfarlane. Reviews. ICES Journal of Marine Science, 18(2):254–255, August 1952. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/18/2/254/887492.

Mackereth:1959:R

[Mac59] F. J. H. Mackereth. Reviews. ICES Journal of Marine Science, 25(1):105–107, October 1959. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/25/1/105/620805.

Margetts:1953:CB

[Mar53] A. R. Margetts. Current bibliography. ICES Journal of Marine Science, 19(2):224–240, October 1953. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/19/2/224/647156.

Margetts:1954:LGR

[Mar54a] A. R. Margetts. The length-girth relationships in haddock and whiting and their application to mesh selection. ICES Journal of Marine Science, 20(1):56–61, July 1954. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/20/1/56/664134.

Marshall:1954:R

[Mar54b] N. B. Marshall. Reviews. ICES Journal of Marine Science, 19(3): 384–386, March 1954. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/19/3/384/662819.

Margetts:1955:CB

[Mar55a] A. R. Margetts. Current bibliography. ICES Journal of Marine Science, 20(3):333–361, May 1955. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/20/3/333/697688.

Margetts:1955:SSM

[Mar55b] A. R. Margetts. Selection of soles by the mesh of trawls. ICES Journal of Marine Science, 20(3):276–289, May 1955. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL

http://academic.oup.com/icesjms/article/20/3/276/697579.

Marr:1956:CPE

[Mar56] John C. Marr. The “critical period” in the early life history of marine fishes. ICES Journal of Marine Science, 21(2):160–170, April 1956. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/21/
2/160/782865.

Margetts:1957:LGR

[Mar57] A. R. Margetts. The length-girth relationships in whiting and cod and their application to mesh selection. ICES Journal of Marine Science, 23(1):64–71, October 1957. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/23/1/64/801800.

Marshall:1958:PPA

[Mar58] P. T. Marshall. Primary production in the Arctic. ICES Journal of Marine Science, 23(2):173–177, April 1958. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/23/2/173/688807.

Marshall:1959:R

[Mar59] N. B. Marshall. Reviews. ICES Journal of Marine Science, 25(1): 102–103, October 1959. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/25/1/102/620799.

Maurin:1952:CMC

[Mau52] Claude Maurin. C. Les merlus des côtes atlantiques du Maroc. (French) [C. The hakes of the Atlantic Coast of Morocco]. ICES Journal of Marine Science, 18(2):224–229, August 1952. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL

http://academic.oup.com/icesjms/article/18/2/224/887461.

Maurin:1954:ECD

[Mau54] Claude Maurin. Étude comparative du “Merlu Blanc” (Merluccius merluccius L.) et du “Merlu Noir” (Merluccius senegalensis Cadenat). (French) [Comparative study of “white hake” (Merluccius merluccius L.) and “black hake” (Merluccius senegalensis Cadenat)]. ICES Journal of Marine Science, 19(3):345–349, March 1954. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/19/
3/345/662744.

Mauchline:1958:IIM

[Mau58] J. Mauchline. An improved instrument for measuring small crustacea. ICES Journal of Marine Science, 23(2):245–248, April 1958. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/23/
2/245/688837.

McIntyre:1953:NMS

[McI53] A. D. McIntyre. A note on market sampling of halibut. ICES Journal of Marine Science, 19(1):69–71, May 1953. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/19/1/69/678898.

Mcl:1953:R

[Mcl53] A. D. Mcl. Reviews. ICES Journal of Marine Science, 18(3):363– 364, January 1953. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/18/3/363/695497.

Mcl:1956:R

[Mcl56] A. D. Mcl. Reviews. ICES Journal of Marine Science, 21(2): 196–197, April 1956. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/21/2/196/782904.

Menon:1950:UBO

[Men50] M. D. Menon. The use of bones, other than otoliths, in determining the age and growth-rate of fishes. ICES Journal of Marine Science, 16(3):311–335, July 1950. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/16/3/311/607800.

Menache:1954:STT

[Men54] M. Menaché. Simplification de la table de titration de M. Knudsen pour la détermination de la chlorinité de l’eau de mer. (French) [Simplification of the titration table of Mr. Knudsen for the determination of the chlorine in sea water]. ICES Journal of Marine Science, 19(3):301–308, March 1954. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/19/3/301/662738.

Menzel:1959:UAG

[Men59] David W. Menzel. Utilization of algae for growth by the angelfish, Holacanthus bermudensis. ICES Journal of Marine Science, 24(2): 308–313, March 1959. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/24/2/308/745041.

Moore:1954:PIL

[MM54] W. H. Moore and C. H. Mortimer. A portable instrument for the location of subcutaneous fish-tags. ICES Journal of Marine Science, 20(1):83–86, July 1954. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/20/1/83/664146.

Molander:1955:SIP

[Mol55] Arvid R. Molander. Swedish investigations on plaice and flounder in the Southern Baltic. ICES Journal of Marine Science, 21(1): 25–43, October 1955. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/21/1/25/630404.

Molander:1957:Ra

[Mol57a] Arvid R. Molander. Reviews. ICES Journal of Marine Science, 22(2):225–226, February 1957. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/22/2/225/612677.

Molander:1957:Rb

[Mol57b] Arvid R. Molander. Reviews. ICES Journal of Marine Science, 22 (3):346–347, June 1957. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/22/3/346/615301.

Morris:1955:SCR

[Mor55] Robert W. Morris. Some considerations regarding the nutrition of marine fish larvae. ICES Journal of Marine Science, 20(3):255–265, May 1955. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/
20/3/255/697570.

Mosby:1950:Ra

[Mos50a] Håkon Mosby. Reviews. ICES Journal of Marine Science, 16(3):388, July 1950. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/
16/3/388/607826.

Mosby:1950:Rb

[Mos50b] Håkon Mosby. Reviews. ICES Journal of Marine Science, 16(3):389, July 1950. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/
16/3/389/607830.

Mosher:1954:UOD

[Mos54] Kenneth H. Mosher. Use of otoliths for determining the age of several fishes from the Bering Sea. ICES Journal of Marine Science, 19(3):337–344, March 1954. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/19/3/337/662743.

Mosby:1958:BHH

[Mos58] Håkon Mosby. Bjørn Helland-Hansen 1877–1957. ICES Journal of Marine Science, 23(3):321–323, July 1958. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/23/3/321/648771.

Muzinic:1958:ARH

[MR58] R. Muzinic and I. D. Richardson. On the appearance of rings on herring scales. ICES Journal of Marine Science, 24(1):120– 134, November 1958. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/24/1/120/721964.

Nedelec:1954:DBT

[Néd54] C. Nédélec. Données biométriques sur le thon rouge de la mer du nord. (French) [Biometric data on North Sea bluefin tuna]. ICES Journal of Marine Science, 19(3):350–355, March 1954. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL

http://academic.oup.com/icesjms/article/19/3/350/662746.

Nicholson:1958:MPP

[Nic58] H. F. Nicholson. Mechanical pipette for picking out diatoms. ICES Journal of Marine Science, 23(2):189–191, April 1958. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL

http://academic.oup.com/icesjms/article/23/2/189/688813.

Nielsen:1950:AAP

[Nie50] Per Haller Nielsen. An auxiliary apparatus for plankton studies by means of the sedimentation method. ICES Journal of Marine Science, 16(3):307–309, July 1950. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/16/3/307/607790.

Nielsen:1952:URA

[Nie52] E. Steemann Nielsen. The use of radio-active carbon (C^{14}) for measuring organic production in the sea. ICES Journal of Marine Science, 18(2):117–140, August 1952. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/18/2/117/887431.

Nielsen:1954:OPO

[Nie54] E. Steemann Nielsen. On organic production in the oceans. ICES Journal of Marine Science, 19(3):309–328, March 1954. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL

http://academic.oup.com/icesjms/article/19/3/309/662739.

Nielsen:1956:R

[Nie56] E. Steemann Nielsen. Reviews. ICES Journal of Marine Science, 22(1):95, October 1956. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/22/1/95/721984.

Nielsen:1958:BBP

[Nie58] E. Steemann Nielsen. The balance between phytoplankton and zooplankton in the sea. ICES Journal of Marine Science, 23(2): 178–188, April 1958. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/23/2/178/688810.

Nielsen:1959:CAD

[Nie59] E. Steemann Nielsen. A central agency for ^{14}C determination (measurements of primary production in the sea) at Charlottenlund Slot, Denmark. ICES Journal of Marine Science, 24(2):372–373, March 1959. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/24/
2/372/745060.

Nikolsky:1958:NRA

[Nik58] G. Nikolsky. Notes on recent Atlantic salmon investigations in the USSR. ICES Journal of Marine Science, 23(3):434–439, July 1958. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/23/
3/434/648830.

Nordli:1950:PSW

[Nor50] Ottar Nordli. Precipitate in sea-water samples for the study of plankton. ICES Journal of Marine Science, 16(3):310, July 1950. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/16/
3/310/607795.

Oppenheimer:1953:WSM

[Opp53] Carl H. Oppenheimer. Why study marine fish diseases? ICES Journal of Marine Science, 19(1):39–43, May 1953. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/19/1/39/678890.

Oren:1952:HL

[Ore52] O. H. Oren. Helmuth Lissner. 1897–1951. ICES Journal of Marine Science, 18(1):7–9, April 1952. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/18/1/7/641197.

Ostvedt:1956:R

[Øst56] O. J. Østvedt. Reviews. ICES Journal of Marine Science, 21(3):331, June 1956. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/
21/3/331/644050.

P:1953:Ra

[P.53a] I. P. Reviews. ICES Journal of Marine Science, 19(1):96–97, May 1953. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/19/
1/96/678934.

P:1953:Rb

[P.53b] I. P. Reviews. ICES Journal of Marine Science, 19(1):97, May 1953. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/19/
1/97/678936.

P:1953:Rc

[P.53c] I. P. Reviews. ICES Journal of Marine Science, 19(1):97, May 1953. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/19/
1/97/678939.

P:1954:Rb

[P.54a] B. B. P. Reviews. ICES Journal of Marine Science, 20(1):95–97, July 1954. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/
20/1/95/664161.

P:1954:Ra

[P.54b] F. T. K. P. Reviews. ICES Journal of Marine Science, 19(3): 390–391, March 1954. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/19/3/390/662840.

P:1955:R

[P.55] J. A. P. Reviews. ICES Journal of Marine Science, 20(3):329–330, May 1955. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/
20/3/329/697676.

P:1956:Ra

[P.56a] J. P. Reviews. ICES Journal of Marine Science, 21(2):193, April 1956. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/21/
2/193/782887.

P:1956:Rb

[P.56b] K. A. P. Reviews. ICES Journal of Marine Science, 22(1):100–101, October 1956. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/
22/1/100/721850.

P:1957:R

[P.57] J. P. Reviews. ICES Journal of Marine Science, 22(2):223– 225, February 1957. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/22/2/223/612670.

Palosuo:1951:IBM

[Pal51] Erkki Palosuo. Ice in the Baltic and the meteorological factors. ICES Journal of Marine Science, 17(2):124–132, April 1951. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/17/
2/124/555558.

Parrish:1951:FCL

[Par51] B. B. Parrish. Fishing capacities of Lowestoft and Aberdeen trawls when used on flatfish grounds. ICES Journal of Marine Science, 17 (2):156–171, April 1951. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/17/2/156/555595.

Parrish:1952:R

[Par52] B. B. Parrish. Reviews. ICES Journal of Marine Science, 18(2): 251–252, August 1952. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/18/2/251/887489.

Parrish:1956:R

[Par56] B. B. Parrish. Reviews. ICES Journal of Marine Science, 21(2): 195–196, April 1956. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/21/2/195/782896.

Parrish:1959:R

[Par59] B. B. Parrish. Reviews. ICES Journal of Marine Science, 24(2): 376–378, March 1959. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/24/2/376/745067.

Pentelow:1950:R

[Pen50] F. T. K. Pentelow. Reviews. ICES Journal of Marine Science, 17(1): 72–73, October 1950. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/17/1/72/627967.

Pentelow:1954:R

[Pen54] F. T. K. Pentelow. Reviews. ICES Journal of Marine Science, 19(3):391–392, March 1954. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/19/3/391/662842.

Perryman:1950:SOC

[Per50] A. D. Perryman. A simple otolith cutter. ICES Journal of Marine Science, 17(1):37–39, October 1950. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/17/1/37/627935.

Paquette:1957:ICB

[PF57] Robert G. Paquette and Herbert F. Frolander. Improvements in the Clarke-bumpus plankton sampler. ICES Journal of Marine Science, 22(3):284–288, June 1957. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/22/3/284/615242.

Pickford:1950:R

[Pic50] G. E. Pickford. Reviews. ICES Journal of Marine Science, 17(1): 66–67, October 1950. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/17/1/66/627952.

Picotti:1958:NTB

[Pic58] Mario Picotti. Nouveaux types de bouteilles à eau en verre et leur emploi en océanographie, en limnologie et en bactériologie. (French) [New types of glass water bottles and their use in oceanography, limnology and bacteriology]. ICES Journal of Marine Science, 24 (1):3–9, November 1958. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/24/1/3/722039.

Parrish:1956:CMM

[PJP56] B. B. Parrish, R. Jones, and J. A. Pope. A comparison of meshmeasuring methods. ICES Journal of Marine Science, 21(3):310– 318, June 1956. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/21/3/310/644031.

Poole:1954:R

[Poo54] H. H. Poole. Reviews. ICES Journal of Marine Science, 19(3): 375–376, March 1954. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/19/3/375/662770.

Postel:1954:CET

[Pos54] E. Postel. Contribution à l’Étude des Thonidés de l’Atlantique tropical. (French) [Contribution to the study of tuna in the Tropical Atlantic]. ICES Journal of Marine Science, 19(3):356–362, March 1954. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/19/
3/356/662747.

Poulsen:1950:R

[Pou50] Erik M. Poulsen. Reviews. ICES Journal of Marine Science, 16 (2):249, January 1950. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/16/2/249/668608.

Poulsen:1951:Ra

[Pou51a] Erik M. Poulsen. Reviews. ICES Journal of Marine Science, 17(3): 318–319, August 1951. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/17/3/318/628699.

Poulsen:1951:Rb

[Pou51b] Erik M. Poulsen. Reviews. ICES Journal of Marine Science, 17(3): 319–320, August 1951. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/17/3/319/628706.

Poulsen:1952:HB

[Pou52] Erik M. Poulsen. Harald Blegvad. 1886–1951. ICES Journal of Marine Science, 18(1):3–6, April 1952. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/18/1/3/641181.

Proudman:1950:R

[Pro50] J. Proudman. Reviews. ICES Journal of Marine Science, 17(1): 63–64, October 1950. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/17/1/63/627948.

Parrish:1959:OTA

[PS59] B. B. Parrish and D. P. Sharman. Otolith types amongst summer– autumn spawning herring in the Northern North Sea. ICES Journal of Marine Science, 25(1):81–92, October 1959. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/25/1/81/620821.

Qasim:1956:TDS

[Qas56] S. Z. Qasim. Time and duration of the spawning season in some marine teleosts in relation to their distribution. ICES Journal of Marine Science, 21(2):144–155, April 1956. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/21/2/144/782853.

R:1950:R

[R.50] F. S. R. Reviews. ICES Journal of Marine Science, 16(3):393, July 1950. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/
16/3/393/607850.

R:1952:R

[R.52] I. R. Reviews. ICES Journal of Marine Science, 18(1):80, April 1952. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/18/
1/80/641217.

R:1953:R

[R.53] B. B. R. Reviews. ICES Journal of Marine Science, 18(3):366–367, January 1953. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/
18/3/366/695508.

R:1954:R

[R.54] W. E. R. Reviews. ICES Journal of Marine Science, 19(3):386–387, March 1954. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/
19/3/386/662823.

R:1955:Ra

[R.55a] K. M. R. Reviews. ICES Journal of Marine Science, 20(3):318–319, May 1955. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/
20/3/318/697636.

R:1955:Rb

[R.55b] K. M. R. Reviews. ICES Journal of Marine Science, 20(3):319–320, May 1955. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/
20/3/319/697641.

R:1956:R

[R.56] K. M. R. Reviews. ICES Journal of Marine Science, 21(2):193–195, April 1956. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/
21/2/193/782892.

Rae:1951:CPW

[Rae51] Bennet B. Rae. A comparison of the pre-war and post-war stocks of lemon soles at Faroe. ICES Journal of Marine Science, 17(3): 242–260, August 1951. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/17/3/242/628680.

Rass:1959:BFC

[Ras59] T. S. Rass. Biogeographical fishery complexes of the Atlantic and Pacific Oceans and their comparison. ICES Journal of Marine Science, 24(2):243–254, March 1959. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/24/2/243/745000.

Rees:1950:R

[Ree50] W. J. Rees. Reviews. ICES Journal of Marine Science, 17(1): 64–65, October 1950. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/17/1/64/627951.

Rees:1953:R

[Ree53] C. B. Rees. Reviews. ICES Journal of Marine Science, 19(2): 222–223, October 1953. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/19/2/222/647152.

Reynolds:1959:FEM

[Rey59] N. Reynolds. Further experiments on mussel purification. ICES Journal of Marine Science, 24(3):486–493, June 1959. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL

http://academic.oup.com/icesjms/article/24/3/486/677832.

Richardson:1958:DMU

[RG58] I. D. Richardson and J. A. Gulland. Description of model used to demonstrate dynamics of exploited fish stocks. ICES Journal of Marine Science, 24(1):55–61, November 1958. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/24/1/55/722069.

Richardson:1954:R

[Ric54] I. D. Richardson. Reviews. ICES Journal of Marine Science, 19(3): 389–390, March 1954. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/19/3/389/662836.

Riley:1953:LE

[Ril53] Gordon A. Riley. Letter to the Editor. ICES Journal of Marine Science, 19(1):85–89, May 1953. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/19/1/85/678906.

Roberts:1950:PEC

[Rob50] B. A. C. H. Roberts. A photo-electric cell comparator for colorimetric analyses at sea. ICES Journal of Marine Science, 17(1): 17–24, October 1950. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/17/1/17/627927.

Roberts:1952:ACW

[Rob52] C. H. Roberts. An automatic collector for water and plankton samples. ICES Journal of Marine Science, 18(2):107–116, August 1952. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/18/
2/107/887426.

Rochford:1956:R

[Roc56] D. J. Rochford. Reviews. ICES Journal of Marine Science, 21 (3):328–331, June 1956. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/21/3/328/644048.

Rollefsen:1953:SDF

[Rol53] Gunnar Rollefsen. The selectivity of different fishing gear used in Lofoten. ICES Journal of Marine Science, 19(2):191–194, October 1953. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/
19/2/191/647073.

Rosen:1952:R

[Ros52] Nils Rosén. Reviews. ICES Journal of Marine Science, 18(1):77–78, April 1952. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/
18/1/77/641212.

Rossiter:1957:R

[Ros57] J. R. Rossiter. Reviews. ICES Journal of Marine Science, 22(3): 341–343, June 1957. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/22/3/341/615281.

Rouch:1951:R

[Rou51] J. Rouch. Reviews. ICES Journal of Marine Science, 17(2):197–198, April 1951. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/
17/2/197/555612.

Russell:1951:RSC

[Rus51] E. S. Russell. Robert Selbie Clark. 1882–1950. ICES Journal of Marine Science, 17(2):99–102, April 1951. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/17/2/99/555637.

Russell:1952:PJ

[Rus52] F. S. Russell. Poul Jespersen. 1891–1951. ICES Journal of Marine Science, 18(2):104–106, August 1952. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/18/2/104/887421.

Ruud:1955:R

[Ruu55] Johan T. Ruud. Reviews. ICES Journal of Marine Science, 20 (3):326–327, May 1955. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/20/3/326/697666.

Ryther:1954:COM

[RV54] J. H. Ryther and R. F. Vaccaro. A comparison of the oxygen and 14C methods of measuring marine photosynthesis. ICES Journal of Marine Science, 20(1):25–34, July 1954. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/20/1/25/664126.

S:1951:R

[S.51] A. S. Reviews. ICES Journal of Marine Science, 17(2):201–202, April 1951. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/
17/2/201/555630.

S:1955:R

[S.55] G. A. S. Reviews. ICES Journal of Marine Science, 20(3):326, May 1955. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/
20/3/326/697662.

S:1957:Ra

[S.57a] J. S. Reviews. ICES Journal of Marine Science, 22(2):212– 213, February 1957. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/22/2/212/612614.

S:1957:Rc

[S.57b] M. S. Reviews. ICES Journal of Marine Science, 22(2):215– 219, February 1957. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/22/2/215/612644.

S:1957:Rb

[S.57c] O. H. S. Reviews. ICES Journal of Marine Science, 22(2):214– 215, February 1957. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/22/2/214/612634.

S:1958:Ra

[S.58a] O. H. S. Reviews. ICES Journal of Marine Science, 23(2):258–259, April 1958. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/
23/2/258/688850.

S:1958:Rb

[S.58b] O. H. S. Reviews. ICES Journal of Marine Science, 23(2):259, April 1958. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/
23/2/259/688854.

S:1958:Rd

[S.58c] S. S. Reviews. ICES Journal of Marine Science, 23(2):269–270, April 1958. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/
23/2/269/688903.

S:1958:Rc

[S.58d] U. S. Reviews. ICES Journal of Marine Science, 23(2):261–262, April 1958. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/
23/2/261/688865.

Strickland:1959:DEA

[SA59] J. D. H. Strickland and K. H. Austin. The direct estimation of ammonia in sea water with notes on “reactive” iron, nitrate, and inorganic phosphorus. ICES Journal of Marine Science, 24(3):446–451, June 1959. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/
24/3/446/677820.

Saelen:1956:Ra

[Sæl56a] O. H. Sælen. Reviews. ICES Journal of Marine Science, 21(2):192, April 1956. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/
21/2/192/782881.

Saelen:1956:Rb

[Sæl56b] O. H. Sælen. Reviews. ICES Journal of Marine Science, 21(3): 327–328, June 1956. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/21/3/327/644047.

Saelert:1956:R

[Sæl56c] O. H. Sælert. Reviews. ICES Journal of Marine Science, 21(2): 192–193, April 1956. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/21/2/192/782884.

Saetersdal:1956:R

[Sæt56] G. Sætersdal. Reviews. ICES Journal of Marine Science, 22(1): 97–98, October 1956. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/22/1/97/722000.

Saetersdal:1958:R

[Sæt58] G. Sætersdal. Reviews. ICES Journal of Marine Science, 23(3): 463–465, July 1958. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/23/3/463/648848.

Saetersdal:1959:R

[Sæt59] G. Sætersdal. Reviews. ICES Journal of Marine Science, 24(3): 502–503, June 1959. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/24/3/502/677844.

Salzen:1957:TSG

[Sal57] E. A. Salzen. A trawling survey off the gold Coast. ICES Journal of Marine Science, 23(1):72–82, October 1957. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/23/1/72/801814.

SanchezySanchez:1957:CON

[San57] Manuel Sanchez y Sanchez. The culture of oysters in North-Western Spain. ICES Journal of Marine Science, 22(2):197–199, February 1957. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/22/
2/197/612597.

Saville:1958:MSP

[Sav58] A. Saville. Mesh selection in plankton nets. ICES Journal of Marine Science, 23(2):192–201, April 1958. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/23/2/192/688816.

Schleicher:1956:CBM

[SB56] Karl E. Schleicher and Alvin Bradshaw. A conductivity bridge for measurement of the salinity of sea water. ICES Journal of Marine Science, 22(1):9–20, October 1956. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/22/1/9/721948.

Schell:1952:RII

[Sch52] I. I. Schell. On the role of the ice off Iceland in the decadal air temperatures of Iceland, and some other areas. ICES Journal of Marine Science, 18(1):11–36, April 1952. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/18/1/11/641179.

Segerstraale:1951:RIS

[Seg51] Sven G. Segerstråle. The recent increase in salinity off the coasts of Finland and its influence upon the fauna. ICES Journal of Marine Science, 17(2):103–110, April 1951. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/17/2/103/555542.

Segerstraale:1956:R

[Seg56] S. Segerstråle. Reviews. ICES Journal of Marine Science, 21(3): 332–333, June 1956. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/21/3/332/644051.

Simpson:1952:WBS

[SG52] A. C. Simpson and Michael Graham. Winds and brood strengths. ICES Journal of Marine Science, 18(1):73, April 1952. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL

http://academic.oup.com/icesjms/article/18/1/73/641199.

Simpsion:1950:CB

[Sim50a] A. C. Simpsion. Current bibliography. ICES Journal of Marine Science, 17(1):78–95, October 1950. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/17/1/78/628042.

Simpson:1950:R

[Sim50b] A. C. Simpson. Reviews. ICES Journal of Marine Science, 17(1): 71–72, October 1950. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/17/1/71/627965.

Simojoki:1951:CSW

[Sim51a] Heikki Simojoki. On the cooling of sea water and the influence of the water heat content on the time of freezing in the Archipelago and North Baltic. ICES Journal of Marine Science, 17(2):121–123, April 1951. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/
17/2/121/555548.

Simpson:1951:R

[Sim51b] A. C. Simpson. Reviews. ICES Journal of Marine Science, 17(2): 198–199, April 1951. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/17/2/198/555615.

Simpson:1953:CB

[Sim53a] A. C. Simpson. Current bibliography. ICES Journal of Marine Science, 19(1):108–123, May 1953. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/19/1/108/678882.

Simpson:1953:SOM

[Sim53b] A. C. Simpson. Some observations on the mortality of fish and the distribution of plankton in the Southern North Sea during the Cold Winter, 1946–1947. ICES Journal of Marine Science, 19(2): 150–177, October 1953. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/19/2/150/647054.

Simpson:1954:R

[Sim54] A. C. Simpson. Reviews. ICES Journal of Marine Science, 20 (1):97–99, July 1954. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/20/1/97/664167.

Simpson:1955:R

[Sim55] A. C. Simpson. Reviews. ICES Journal of Marine Science, 21 (1):80, October 1955. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/21/1/80/630448.

Sjoblom:1958:CPC

[Sjö58] Veikko Sjöblom. On the causes of the poor catches of Baltic herring on the Finnish Coast in the summer of 1956. ICES Journal of Marine Science, 24(1):111–119, November 1958. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL

http://academic.oup.com/icesjms/article/24/1/111/721959.

Smayda:1959:SIR

[Sma59] Theodore J. Smayda. The seasonal incoming radiation in Norwegian and Arctic waters, and indirect methods of measurement. ICES Journal of Marine Science, 24(2):215–220, March 1959. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL

http://academic.oup.com/icesjms/article/24/2/215/744984.

Somme:1950:R

[Søm50] Sven Sømme. Reviews. ICES Journal of Marine Science, 17(1): 73–74, October 1950. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/17/1/73/627986.

Somme:1951:EJD

[Søm51] Sven Sømme. The Editor journal du conseil. ICES Journal of Marine Science, 17(3):316, August 1951. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/17/3/316/628690.

Somme:1952:KD

[Søm52] Sven Sømme. Knut Dahl. 1871–1951. ICES Journal of Marine Science, 18(2):101–103, August 1952. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/18/2/101/887420.

Somme:1958:Ra

[Søm58a] O. M. Sømme. Reviews. ICES Journal of Marine Science, 24 (1):157–158, November 1958. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/24/1/157/722006.

Somme:1958:Rb

[Søm58b] O. M. Sømme. Reviews. ICES Journal of Marine Science, 24 (1):158–160, November 1958. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/24/1/158/722012.

Souter:1950:HRM

[Sou50] W. C. Souter. Hugh Robert Mill. 1861–1950. ICES Journal of Marine Science, 17(1):13–15, October 1950. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/17/1/13/627926.

Steers:1954:R

[Ste54] J. A. Steers. Reviews. ICES Journal of Marine Science, 19(3): 369–370, March 1954. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/19/3/369/662755.

Steele:1957:RLE

[Ste57a] John H. Steele. The role of lateral eddy diffusion in the Northern North Sea. ICES Journal of Marine Science, 22(2):152–162, February 1957. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/
22/2/152/612557.

Steven:1957:R

[Ste57b] G. A. Steven. Reviews. ICES Journal of Marine Science, 23(1):103, October 1957. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/
23/1/103/801705.

Stott:1958:R

[Sto58] B. Stott. Reviews. ICES Journal of Marine Science, 23(2):268, April 1958. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/
23/2/268/688898.

Strasburg:1959:INM

[Str59] Donald W. Strasburg. An instance of natural mass mortality of larval frigate mackerel in the Hawaiian Islands. ICES Journal of Marine Science, 24(2):255–263, March 1959. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/24/2/255/745004.

Sundnes:1957:TLC

[Sun57] Gunnar Sundnes. On the transport of live cod and coalfish. ICES Journal of Marine Science, 22(2):191–196, February 1957. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL

http://academic.oup.com/icesjms/article/22/2/191/612594.

Svardson:1952:R

[Svä52] G. Svärdson. Reviews. ICES Journal of Marine Science, 18(1): 79–80, April 1952. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/18/1/79/641214.

Sverdrup:1953:CVB

[Sve53] H. U. Sverdrup. On conditions for the vernal blooming of phytoplankton. ICES Journal of Marine Science, 18(3):287–295, January 1953. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/
18/3/287/695418.

Swain:1957:R

[Swa57] A. Swain. Reviews. ICES Journal of Marine Science, 22(3):348–349, June 1957. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/
22/3/348/615308.

T:1950:R

[T.50] H. T. Reviews. ICES Journal of Marine Science, 16(3):388–389, July 1950. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/
16/3/388/607828.

T:1957:R

[T.57] H. J. T. Reviews. ICES Journal of Marine Science, 22(3):351–352, June 1957. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/
22/3/351/615322.

Tait:1950:R

[Tai50] J. B. Tait. Reviews. ICES Journal of Marine Science, 17(1):62, October 1950. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/
17/1/62/627947.

Tait:1951:AMR

[Tai51] John B. Tait. On the application to marine research of the meteorological concept of accumulated temperature. ICES Journal of Marine Science, 17(2):111–120, April 1951. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/17/2/111/555545.

Tait:1958:HT

[Tai58] J. B. Tait. Harold Thompson 1890–1957. ICES Journal of Marine Science, 23(2):151–154, April 1958. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/23/2/151/688778.

Taylor:1958:CGT

[Tay58] Clyde C. Taylor. Cod growth and temperature. ICES Journal of Marine Science, 23(3):366–370, July 1958. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/23/3/366/648803.

Taylor:1959:TGP

[Tay59] Clyde C. Taylor. Temperature and growth — the Pacific razor clam. ICES Journal of Marine Science, 25(1):93–101, October 1959. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/25/
1/93/620823.

Tesch:1950:FL

[Tes50] J. J. Tesch. François Liebert. 1882–1918. ICES Journal of Marine Science, 16(2):164–165, January 1950. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/16/2/164/668462.

Truesdale:1957:SOS

[TG57] G. A. Truesdale and A. L. H. Gameson. The solubility of oxygen in saline water. ICES Journal of Marine Science, 22(2):163– 166, February 1957. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/22/2/163/612565.

Th:1957:R

[Th.57] H. Th. Reviews. ICES Journal of Marine Science, 22(2):213– 214, February 1957. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/22/2/213/612618.

Thomsen:1950:MK

[Tho50] Helge Thomsen. Martin Knudsen. 1871–1949. ICES Journal of Marine Science, 16(2):155–159, January 1950. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/16/2/155/668453.

Thompson:1952:CSH

[Tho52] W. F. Thompson. Condition of stocks of halibut in the Pacific. ICES Journal of Marine Science, 18(2):141–166, August 1952. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL

http://academic.oup.com/icesjms/article/18/2/141/887435.

Thomas:1953:EFM

[Tho53] H. J. Thomas. The efficiency of fishing methods employed in the capture of lobsters and crabs. ICES Journal of Marine Science, 18(3):333–350, January 1953. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/18/3/333/695444.

Thomas:1954:ECP

[Tho54] H. J. Thomas. The efficiency of the Cornish pot and the Scottish creel in the capture of lobsters and crabs. ICES Journal of Marine Science, 20(1):87–91, July 1954. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/20/1/87/664149.

Thomas:1955:OSR

[Tho55] H. J. Thomas. Observations on the sex ratio and mortality rates in the lobster (Homarus vulgaris edw.). ICES Journal of Marine Science, 20(3):295–305, May 1955. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/20/3/295/697585.

Thomas:1958:SSV

[Tho58] H. J. Thomas. Some seasonal variations in the catch and stock composition of the lobster. ICES Journal of Marine Science, 24 (1):147–154, November 1958. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/24/1/147/721981.

Thomas:1959:CSC

[Tho59] H. J. Thomas. Creel selectivity in the capture of lobsters and crabs. ICES Journal of Marine Science, 24(2):342–348, March 1959. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/24/
2/342/745046.

Truesdale:1956:SRW

[TK56] G. A. Truesdale and G. Knowles. Some recent work on dissolved oxygen in natural waters. ICES Journal of Marine Science, 21 (3):263–267, June 1956. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/21/3/263/644016.

Tomczak:1954:R

[Tom54] G. Tomczak. Reviews. ICES Journal of Marine Science, 20(2):210– 211, December 1954. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/20/2/210/780556.

Trewavas:1950:R

[Tre50] E. Trewavas. Reviews. ICES Journal of Marine Science, 16(2): 254–257, January 1950. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/16/2/254/668660.

Trewavas:1954:R

[Tre54] E. Trewavas. Reviews. ICES Journal of Marine Science, 20(2):216– 217, December 1954. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/20/2/216/780599.

Trout:1958:REC

[Tro58] Geoffrey C. Trout. Results of English cod tagging in the Barents Sea. ICES Journal of Marine Science, 23(3):371–380, July 1958. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/23/
3/371/648812.

Tucker:1951:R

[Tuc51] Denys W. Tucker. Reviews. ICES Journal of Marine Science, 17 (2):199–200, April 1951. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/17/2/199/555620.

Tucker:1958:R

[Tuc58] M. J. Tucker. Reviews. ICES Journal of Marine Science, 23(2): 259–261, April 1958. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/23/2/259/688860.

Tully:1952:R

[Tul52] John P. Tully. Reviews. ICES Journal of Marine Science, 18(2): 247–249, August 1952. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/18/2/247/887481.

Twomey:1957:SRS

[Two57] Eileen Twomey. Salmon of the River Shannon (1952–1956). ICES Journal of Marine Science, 23(1):89–96, October 1957. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL

http://academic.oup.com/icesjms/article/23/1/89/801825.

Ursin:1954:R

[Urs54] Erik Ursin. Reviews. ICES Journal of Marine Science, 19(3): 382–383, March 1954. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/19/3/382/662810.

Ursell:1957:R

[Urs57] F. Ursell. Reviews. ICES Journal of Marine Science, 22(2):211– 212, February 1957. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/22/2/211/612607.

Vaux:1952:R

[Vau52] David Vaux. Reviews. ICES Journal of Marine Science, 18(2): 246–247, August 1952. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/18/2/246/887478.

Vaux:1953:R

[Vau53] David Vaux. Reviews. ICES Journal of Marine Science, 19(1):95– 96, May 1953. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/
19/1/95/678932.

Vaux:1954:R

[Vau54] David Vaux. Reviews. ICES Journal of Marine Science, 19(3): 376–377, March 1954. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/19/3/376/662776.

vonBrandt:1955:EDN

[vB55] A. von Brandt. The efficiency of drift-nets. ICES Journal of Marine Science, 21(1):8–16, October 1955. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/21/1/8/630444.

Verwey:1950:R

[Ver50] J. Verwey. Reviews. ICES Journal of Marine Science, 17(1):67–68, October 1950. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/
17/1/67/627954.

Vibert:1955:R

[Vib55] R. Vibert. Reviews. ICES Journal of Marine Science, 20(3): 328–329, May 1955. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/20/3/328/697670.

vanLandingham:1957:MKM

[vL57] John W. van Landingham. A modification of the Knudsen method for salinity determination. ICES Journal of Marine Science, 22(2): 174–179, February 1957. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/22/2/174/612579.

Vaccaro:1954:BES

[VR54] R. F. Vaccaro and J. H. Ryther. The bactericidal effects of sunlight in relation to “light” and “dark” bottle photosynthesis experiments. ICES Journal of Marine Science, 20(1):18–24, July 1954. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL

http://academic.oup.com/icesjms/article/20/1/18/664124.

W:1951:Rb

[W.51a] A. E. J. W. Reviews. ICES Journal of Marine Science, 17(3):324, August 1951. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/
17/3/324/628712.

W:1951:Ra

[W.51b] R. S. W. Reviews. ICES Journal of Marine Science, 17(2):202–203, April 1951. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/
17/2/202/555631.

W:1953:Ra

[W.53a] A. E. J. W. Reviews. ICES Journal of Marine Science, 18(3):368, January 1953. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/
18/3/368/695536.

W:1953:Rc

[W.53b] G. D. W. Reviews. ICES Journal of Marine Science, 19(1):106–107, May 1953. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/
19/1/106/678879.

W:1953:Rb

[W.53c] T. W. Reviews. ICES Journal of Marine Science, 19(1):100–102, May 1953. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/
19/1/100/678861.

W:1954:Ra

[W.54a] A. E. J. W. Reviews. ICES Journal of Marine Science, 19(3):392, March 1954. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/
19/3/392/662848.

W:1954:Rb

[W.54b] G. D. W. Reviews. ICES Journal of Marine Science, 20(1):100–101, July 1954. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/
20/1/100/664111.

W:1955:R

[W.55] Kr. Fr. W. Reviews. ICES Journal of Marine Science, 21(1):78–80, October 1955. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/
21/1/78/630440.

W:1956:R

[W.56] G. D. W. Reviews. ICES Journal of Marine Science, 22(1):104–105, October 1956. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/
22/1/104/721866.

Walker:1954:DLA

[Wal54] F. T. Walker. Distribution of laminariaceae around Scotland. ICES Journal of Marine Science, 20(2):160–166, December 1954. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL

http://academic.oup.com/icesjms/article/20/2/160/780495.

Walne:1956:DCO

[Wal56] P. R. Walne. Destruction of competitive organisms on artificial oyster-spat collectors. ICES Journal of Marine Science, 22(1):75– 76, October 1956. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/22/1/75/721932.

Warren:1958:DIR

[War58a] Peter J. Warren. Description of an improved rotating counter for zooplankton. ICES Journal of Marine Science, 23(3):337–339, July 1958. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/23/
3/337/648780.

Warren:1958:DRS

[War58b] Peter J. Warren. A device for rapid single-handed measurement of shellfish. ICES Journal of Marine Science, 23(3):440–442, July 1958. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/23/
3/440/648835.

Waugh:1957:R

[Wau57] G. D. Waugh. Reviews. ICES Journal of Marine Science, 23(1):107, October 1957. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/
23/1/107/801737.

Weenink:1954:R

[Wee54] M. P. H. Weenink. Reviews. ICES Journal of Marine Science, 19(3):370–371, March 1954. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/19/3/370/662760.

Weenink:1958:R

[Wee58] M. P. H. Weenink. Reviews. ICES Journal of Marine Science, 23 (2):264–266, April 1958. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/23/2/264/688882.

Went:1950:GPF

[Wen50a] Arthur E. J. Went. George Philip Farran. 1876–1949. ICES Journal of Marine Science, 16(2):160–163, January 1950. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL

http://academic.oup.com/icesjms/article/16/2/160/668457.

Went:1950:SRS

[Wen50b] Arthur E. J. Went. Salmon of the River Shannon in 1946 and 1947. ICES Journal of Marine Science, 16(3):341–357, July 1950. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/16/
3/341/607812.

Went:1953:SRS

[Wen53] Arthur E. J. Went. Salmon of the River Shannon (1948–1951). ICES Journal of Marine Science, 18(3):326–332, January 1953. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL

http://academic.oup.com/icesjms/article/18/3/326/695437.

Went:1957:Ra

[Wen57a] A. E. J. Went. Reviews. ICES Journal of Marine Science, 23 (1):105, October 1957. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/23/1/105/801714.

Went:1957:Rb

[Wen57b] A. E. J. Went. Reviews. ICES Journal of Marine Science, 23(1): 106–107, October 1957. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/23/1/106/801722.

Westenberg:1953:AAS

[Wes53a] J. Westenberg. Acoustical aspects of some Indonesian fisheries. ICES Journal of Marine Science, 18(3):311–325, January 1953. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/18/
3/311/695433.

Westenberg:1953:R

[Wes53b] J. Westenberg. Reviews. ICES Journal of Marine Science, 19(1): 102–103, May 1953. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/19/1/102/678863.

Westenberg:1955:R

[Wes55] J. Westenberg. Reviews. ICES Journal of Marine Science, 20(3): 321–323, May 1955. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/20/3/321/697647.

Whitney:1957:FAA

[Whi57] G. G. Whitney, Jr. Factors affecting the accuracy of thermometric depth determinations. ICES Journal of Marine Science, 22(2):167– 173, February 1957. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/22/2/167/612570.

Wiborg:1953:R

[Wib53] Kr. Fred. Wiborg. Reviews. ICES Journal of Marine Science, 19(2): 216–217, October 1953. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/19/2/216/647134.

Wiborg:1954:Ra

[Wib54a] Kr. Fr. Wiborg. Reviews. ICES Journal of Marine Science, 19(3): 380–381, March 1954. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/19/3/380/662802.

Wiborg:1954:Rb

[Wib54b] Kr. Fr. Wiborg. Reviews. ICES Journal of Marine Science, 19(3): 381–382, March 1954. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/19/3/381/662805.

Wiborg:1955:R

[Wib55] K. F. Wiborg. Reviews. ICES Journal of Marine Science, 20(3): 317–318, May 1955. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/20/3/317/697634.

Wickstead:1958:SLZ

[Wic58] J. H. Wickstead. A survey of the larger zooplankton of Singapore Straits. ICES Journal of Marine Science, 23(3):340–353, July 1958. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/23/
3/340/648784.

Wimpenny:1958:RES

[Wim58] R. S. Wimpenny. Robert Edward Savage 1887–1957. ICES Journal of Marine Science, 23(3):324–326, July 1958. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/23/3/324/648775.

Wise:1958:WSI

[Wis58] John P. Wise. The World’s southernmost indigenous cod. ICES Journal of Marine Science, 23(2):208–212, April 1958. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL

http://academic.oup.com/icesjms/article/23/2/208/688821.

Wesenberg-Lund:1958:R

[WL58] E. Wesenberg-Lund. Reviews. ICES Journal of Marine Science, 23 (3):462–463, July 1958. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/23/3/462/648846.

Wood:1955:FMM

[Woo55] E. J. Ferguson Wood. Fluorescent microscopy in marine microbiology. ICES Journal of Marine Science, 21(1):6–7, October 1955. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/21/
1/6/630412.

Wood:1956:CP

[Woo56] E. J. Ferguson Wood. Considerations on productivity. ICES Journal of Marine Science, 21(3):280–283, June 1956. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/21/3/280/644021.

Wood:1957:FAP

[Woo57] P. C. Wood. Factors affecting the pollution and self-purification of molluscan shellfish. ICES Journal of Marine Science, 22(2):200– 208, February 1957. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/22/2/200/612602.

Wood:1958:FCN

[Woo58] R. J. Wood. Fat cycles of North Sea herring. ICES Journal of Marine Science, 23(3):390–398, July 1958. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/23/3/390/648820.

Wood:1959:IGH

[Woo59] R. J. Wood. Investigations on 0-group herring. ICES Journal of Marine Science, 24(2):264–276, March 1959. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/24/2/264/745011.

Wood:1950:ESE

[WP50] Henry Wood and B. B. Parrish. Echo-sounding experiments on fishing gear in action. ICES Journal of Marine Science, 17(1):25–36, October 1950. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/
17/1/25/627928.

Walker:1957:PCL

[WR57a] F. T. Walker and W. D. Richardson. Perennial changes of Laminaria cloustoni EDM. on the coasts of Scotland. ICES Journal of Marine Science, 22(3):298–308, June 1957. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/22/3/298/615255.

Woodhead:1957:SPM

[WR57b] P. M. J. Woodhead and J. D. Riley. The separation of potential males and females in stage-v copepodites of Calanus helgolandicus. ICES Journal of Marine Science, 23(1):47–50, October 1957. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/
23/1/47/801777.

Woodhead:1959:SSC

[WR59] P. M. J. Woodhead and J. D. Riley. Separation of the sexes of Calanus finmarchicus (Gunn.) in the fifth copepodite stage, with comments on the sex ratio and the duration in this stage. ICES Journal of Marine Science, 24(3):465–471, June 1959. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL

http://academic.oup.com/icesjms/article/24/3/465/677827.

Wulff:1956:EIH

[Wul56] A. Wulff. Über das Eingeweidefett (“Ister”) des Herings. (German) [On the belly fat (“Ister”) of herring]. ICES Journal of Marine Science, 21(2):171–174, April 1956. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/21/2/171/782872.

Wust:1954:R

[Wüs54] Georg Wüst. Reviews. ICES Journal of Marine Science, 20(2):209– 210, December 1954. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/20/2/209/780547.

Wust:1955:Rb

[Wüs55a] G. Wüst. Reviews. ICES Journal of Marine Science, 20(3):315, May 1955. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/
20/3/315/697614.

Wust:1955:Ra

[Wüs55b] Georg Wüst. Reviews. ICES Journal of Marine Science, 20(3): 311–313, May 1955. CODEN ICESEC. ISSN 1054-3139 (print),

1095-9289 (electronic). URL http://academic.oup.com/icesjms/
article/20/3/311/697603.

Wust:1958:R

[Wüs58] G. Wüst. Reviews. ICES Journal of Marine Science, 23(2):263–264, April 1958. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/
23/2/263/688878.

Wyrtki:1956:MCS

[Wyr56] Klaus Wyrtki. Monthly charts of surface salinity in Indonesian and adjacent waters. ICES Journal of Marine Science, 21(3):268–279, June 1956. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/
21/3/268/644019.

Wyrtki:1958:LE

[Wyr58] K. Wyrtki. Letter to the Editor. ICES Journal of Marine Science, 24(1):155, November 1958. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/24/1/155/721991.

Yentsch:1957:GDP

[YH57] Charles S. Yentsch and J. Frank Hebard. A gauge for determining plankton volume by the mercury immersion method. ICES Journal of Marine Science, 22(2):184–190, February 1957. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL

http://academic.oup.com/icesjms/article/22/2/184/612589.

Yentsch:1959:RSN

[YR59] Charles S. Yentsch and John H. Ryther. Relative significance of the net phytoplankton and nanoplankton in the waters of Vineyard Sound. ICES Journal of Marine Science, 24(2):231–238, March 1959. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/24/
2/231/744991.